### v1

# Proper maps of annuli

Abstract.
We study rational proper holomorphic maps of annuli in complex euclidean spaces, that is, domains with $U(n)$ as the automorphism group. By the Hartogs phenomenon and a result of Forstnerič, such maps are always rational and extend to proper maps of balls. We first prove that a proper map of annuli from $n$ dimensions to $N$ dimensions where $N<\binom{n+1}{2}$ is always an affine embedding. This inequality is sharp as the homogeneous map of degree 2 satisfies $N=\binom{n+1}{2}$. Next we find a necessary and sufficient condition for a map to be homogeneous: A proper map of annuli is homogeneous if and only if its general hyperplane rank, the affine dimension of the image of a general hyperplane, is exactly N1N-1. This result is really a result classifying homogeneous proper maps of balls. A homogeneous proper ball map takes spheres centered at the origin to spheres centered at the origin. We show that if a proper ball map has general hyperplane rank N1N-1 and takes one sphere centered at the origin to a sphere centered at the origin, then it is homogeneous. Another corollary of this result is a complete classification of proper maps of annuli from dimension 2 to dimension 3.
Key words and phrases:
rational sphere maps, proper holomorphic mappings, maps of annuli, Segre varieties, affine dimension
2020 Mathematics Subject Classification:
32H35, 32A08, 32H02
The second author was in part supported by Simons Foundation collaboration grant 710294.
1. Introduction
Let $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ denote the unit ball. A well-studied problem in complex analysis is to attempt to classify the proper holomorphic maps $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. If $n=N=1$, then by a classical theorem of Fatou [fatou-1923-fonctions], $f$ is a finite Blaschke product and hence rational. If $n=N>1$, then by a theorem of Alexander [alexander-1977-proper], $f$ is an automorphism of the unit ball. No maps exist if $N<n$ for elementary reasons. Dor and Løw [dor-1990-proper, low-1985-embeddings] show that if $N>n>1$, there exist non-rational maps, even in the case of codimension 1, that is, $N=n+1$. If in the case $n>1$ one assumes some regularity on the boundary, then by a theorem of Forstnerič [forstneric-1989-extending], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$. Faran [faran-1982-maps] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are equivalent to one of four polynomial maps. For maps that are $C^{3}$ up to the boundary, Webster [webster-1979-mapping] showed that for $n\geq 3$, there is no map from $\mathbb{B}_{n}$ to $\mathbb{B}_{n+1}$ that is not equivalent up to automorphisms (spherically equivalent) to the linear embedding $z\oplus 0$. Faran [faran-1986-linearity] proved the same conclusion (only the linear embedding) for maps real-analytic up to the boundary for all $N$ up to N=2n2N=2n-2. Huang [huang-1999-linearity] improved the regularity to $C^{2}$. For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding. The map $z^{\otimes d}$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$. This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary mapping to what we will call $H_{d}$ direct sum a number of zero components. Rudin [rudin-1984-homogeneous] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$. See D’Angelo [dangelo-1988-polynomial] for a simpler proof of this fact. In general, see the books by D’Angelo [dangelo-1993-several, dangelo-2019-hermitian, dangelo-2021-rational] for a more thorough treatment of proper maps of balls.
The ball can be regarded as the strongly pseudoconvex domain with the most symmetry. Let us consider then a non-pseudoconvex domain with a lot of symmetry, the annulus. Write $\mathbb{A}_{n,r}=\mathbb{B}_{n}\setminus\overline{r\mathbb{B}_{n}}=\{z\in\mathbb{C}^{n}:r<\mathopen{}\mathclose{{\left\lVert z}}\right\rVert<1\}$ for an annulus, where we will always assume that $0<r<1$. The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group as long as $n>1$. When $n=1$ there are also the automorphisms $e^{i\theta}\frac{r}{z}$. The same classical argument to classify the automorphism group $\mathbb{A}_{1,r}$ can be extended to show that in one dimension a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case it is either equal to $e^{i\theta}z^{d}$ or $r^{d}e^{i\theta}z^{-d}$. In dimension $n\geq 2$, due to the inner pseudoconcave boundary, via the Hartogs phenomenon and the theorem of Forstnerič on rationality of proper ball maps, all proper holomorphic maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are rational, see Theorem˜2.2, but they can be more complicated. In fact, holomorphic proper maps of annuli are in one to one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero. See [helal-2025-proper] for more details.
For $z\in\mathbb{C}^{n}$, if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
(1) $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
is constant. Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin. Moreover, it is not difficult to show that it takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N$ is the number of components of $H_{d}$ and $R=r^{d}$. Similarly, if we have any two proper maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$. We call such a weighted direct sum a juxtaposition. So juxtapositions of proper maps of annuli are proper maps of annuli as long as the source annulus is the same. Based on the one-dimensional result, noting that $z^{\otimes d}$ is the correct multidimensional analogue of $z^{d}$, one could (incorrectly) conjecture that all such proper maps are juxtapositions of homogeneous maps. However, other maps do exist, see Example˜2 and more generally [helal-2025-proper].
Our goal is to start a classification of proper maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$. We say $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ and $g\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are unitarily equivalent (or simply equivalent) if there exist unitary matrices $V\in U(n)$ and $W\in U(N)$ so that
(2) $f=WgV.$
As we noted, $U(n)$ is the holomorphic automorphism group of $\mathbb{A}_{n,r}$. If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.6. If $N>n$, then an affine embedding exists whenever $r\leq R<1$, but is not necessarily $z\oplus 0$ due to the different automorphism group. In particular, $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
(3) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0,$
where the last 0 has Nn1N-n-1 terms. That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component. See Proposition˜2.7.
As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney mapping exists and is not spherically equivalent to a linear embedding. Curiously, for annulus maps, this gap is much longer and goes until $N=\binom{n+1}{2}$.
Theorem 1.1.
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $2\leq n<N<\binom{n+1}{2}$. Then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$. The bound on $N$ is sharp as the homogeneous map $H_{2}$ takes the annulus $\mathbb{A}_{n,r}$ properly to $\mathbb{A}_{\binom{n+1}{2},r^{2}}$.
As homogeneous maps play an outsize role in the study of proper maps of annuli we next find a property that determines a map to be homogeneous. The framework that we find applies more broadly to proper maps of balls and so we state it first in that setting. First we define a certain invariant for proper mappings that has, in various guises, appeared in the study of ball maps. First let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any holomorphic map. If $X\subset\mathbb{C}^{N}$, write $\operatorname{aff}(X)$ for the affine hull of $X$, that is, the smallest affine subspace of $\mathbb{C}^{N}$ containing $X$. Define
(4) $k_{f}\overset{\text{def}}{=}\max\mathopen{}\mathclose{{\left\{\dim\operatorname{aff}\bigl(f(H\cap U)\bigr):H\subset\mathbb{C}^{n}\text{ is an affine hyperplane}}}\right\}.$
We call $k_{f}$ the general hyperplane rank. One could think of $k_{f}$ as a certain measure of complexity. It is not difficult to see that $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps. While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls has $k_{f}\leq N-1$. The idea that Dor [dor-1990-proper] used to show a proper map is not rational is to, in other words, show that the constructed map has $k_{f}=N$ as any rational ball map has $k_{f}=N-1$. If $k_{f}=N-1$, then we obtain a map of Grassmannians. That technique was used for example by Faran to prove his classification of maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The number $k_{f}$ is related to the $X$-variety that Forstnerič used to show the rationality result. In particular, for a rational map, if the dimension of the generic fiber of the $X$-variety is $\ell$, then $k_{f}=N-\ell-1$. D’Angelo [dangelo-2003-Xvariety] has studied the $X$-variety for rational maps and developed a technique of quickly computing fibers of the $X$-variety as the nullity of a certain matrix $C(w)$. The nullity of $C(w)$ for a generic $w$ is then the $\ell$. A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{B}_{n})$ (or $f(\mathbb{A}_{n,r})$). In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{B}_{n})$. In either case, restricting to $S$ on the target we obtain a proper map between balls or annuli with target dimension $N_{f}$. While for rational proper maps, often $k_{f}=N_{f}-1$, it is not always true. Example˜2 mentioned above has $k_{f}=N_{f}-3$. A juxtaposition of two rational maps also has $k_{f}\leq N_{f}-2$, see Proposition˜2.12. What we prove is that for proper maps of annuli, $k_{f}=N_{f}-1$ is precisely the condition that characterizes homogeneous maps $H_{d}$.
Theorem 1.2.
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
(5) $f=UH_{d},$
for some $d\in\mathbb{N}$ and a unitary $U\in U(N)$.
One direction of this theorem is obvious, it is that $k_{f}=N-1$ implies that $f$ is homogeneous that is remarkable. The proof relies on a result from [helal-2025-proper], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps. We note that if we wish to state the theorem in terms of $N_{f}$, then we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$. We note that $k_{f}=N-1$ implies that $N=N_{f}$ for any proper rational map of annuli (or balls).
Let us state a version of the forward direction of this result for proper maps of balls. Clearly $H_{d}$ is a proper ball map that takes infinitely many spheres centered at the origin to spheres centered at the origin. With the assumption that $k_{f}=N-1$, to show that $f$ is homogeneous, it is sufficient to check that a map takes just one inner sphere centered at the origin to a sphere centered at the origin.
Theorem 1.3.
Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a rational proper map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$. Then there exists a unitary map $U$ and an integer $d$ such that $f=UH_{d}$.
The proof follows by using the $k_{f}=N-1$ condition to show that $f$ must in fact take infinitely many spheres centered at zero to spheres centered at zero.
As a corollary to Theorem˜1.2 and to the result of Faran classifying proper maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$, we completely classify proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$.
Theorem 1.4.
Suppose $f\colon\mathbb{A}_{2,r}\to\mathbb{A}_{3,R}$ is a proper holomorphic map. Then $f$ is unitarily equivalent to exactly one of the following two maps:
  1. (i)
    The affine embedding
    (6) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{1},\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{2},\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}}}\right),$
    where $R\geq r$. Note that $N_{f}=2$ and $k_{f}=1$.
  2. (ii)
    The homogeneous map $H_{2}$
    (7) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(z_{1}^{2},\sqrt{2}\,z_{1}z_{2},z_{2}^{2}}}\right),$
    where $R=r^{2}$. Note that $N_{f}=3$ and $k_{f}=2$.
Let us give the organization of this paper. In Section˜2, we state and prove the basic results for maps of annuli. In Section˜3, we prove the first gap theorem, that is, Theorem˜1.1. In Section˜4, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2. We also prove the classification of annuli maps in dimension 2 and 3, that is, Theorem˜1.4. Finally, in Section˜5, we give some further remarks and open questions.
2. Basic results on proper maps of annuli
Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper mappings of annuli in one dimension.
Proposition 2.1.
Suppose $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ is a proper holomorphic map. Then there is some $d\in\mathbb{N}$ and $\theta\in\mathbb{R}$ so that $R=r^{d}$ and $f(z)=e^{i\theta}z^{d}$ or $f(z)=r^{d}e^{i\theta}z^{-d}$.
Sketch of proof.
Let us sketch the proof as it may not be easy to find, but it is almost identical to classifying biholomorphisms of annuli. First using an inversion and scaling and the maximum principle, we may assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner. The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and goes to 0 on the outer circle and 1 on the inner circle. Hence it is equal to $\frac{\log\mathopen{}\mathclose{{\left\lvert z}}\right\rvert}{\log r}$, and we get $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{\frac{\log R}{\log r}}$. So $f(z)=e^{i\theta}z^{d}$ and $R=r^{d}$ for some $d\in\mathbb{N}$. ∎
In [helal-2025-proper], the authors have proved, using the theorem of Forstnerič [forstneric-1989-extending] as a key ingredient, that all annulus maps are rational when $n\geq 2$. For completeness we reproduce it here.
Theorem 2.2 ([helal-2025-proper]).
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map. Then $f$ is a rational map of degree bounded by some expression $D(n,N)$. Moreover, $f$ extends to a holomorphic map on $\mathbb{B}_{n}$ and gives a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ and also a proper map of $r\mathbb{B}_{n}$ to $R\mathbb{B}_{N}$.
Unlike in the ball case, no hypothesis on the boundary regularity is needed due to the pseudoconcavity of the inner boundary. The best expression for $D$ is not known, though D’Angelo has conjectured [dangelo-2003-sharp] that it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 is $n=2$.
Sketch of proof.
As $n\geq 2$, the map $f$ extends holomorphically through the inner ball via Hartogs phenomenon and the image is contained in $\mathbb{B}_{N}$ by the maximum principle. It must take the inner sphere to the inner sphere as it takes the pseudoconcave part of the boundary of the source to the pseudoconcave part of the boundary of the target. By the theorem of Forstnerič, $f$ restricted to the inner ball $r\mathbb{B}_{n}$ is a proper rational map to $R\mathbb{B}_{N}$ and gives the degree bound. That $f$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ follows since we must take the pseudoconvex part of the boundary to the pseudoconvex part. ∎
Via a theorem of Cima–Suffridge [cima-1990-boundary], a rational proper map of balls has no poles on the spheres, so we get a sphere map. Conversely a rational map taking a sphere to a sphere is a proper ball map via maximum principle. In [helal-2025-proper], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere. Putting all this together one gets the following equivalence.
Corollary 2.3.
Suppose $n\geq 2$. An $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map if and only if $f$ is the restriction of a rational map $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{n}$ with no poles on the spheres $rS^{2n-1}$ nor $S^{2n-1}$ such that $f(rS^{2n-1})\subset RS^{2N-1}$ and $f(S^{2n-1})\subset S^{2N-1}$.
If $n=N=1$ the rationality still holds (but not the bound for the degree) as we just showed above. However, if $N>n=1$, then nonrational maps exist.
Example 1.
Next, let us show that a nonrational proper map exists from $\mathbb{A}_{1,r}$ to $\mathbb{A}_{2,R}$ for some $R$. First, we construct a nonrational proper map from the disc $\mathbb{D}$ to the 2-ball $\mathbb{B}_{2}$. Such a map follows from the result of Dor [dor-1990-proper], but it is useful to consider an elementary construction in the one-dimensional case. Take an arbitrary nonrational function from the disc $f_{1}\colon\mathbb{D}\to\mathbb{D}$, suppose that it extends continuously to the closed disc $\overline{\mathbb{D}}$ but does not extend holomorphically past any point on $\partial\mathbb{D}$ and suppose that $\sup_{z\in\overline{\mathbb{D}}}\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert<1$. Find a real-valued harmonic function $u$ on $\mathbb{D}$ continuous on $\overline{\mathbb{D}}$ so that on $\partial\mathbb{D}$ we have $u(z)=\frac{1}{2}\log\mathopen{}\mathclose{{\left(1-\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}}}\right)$. Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$. Then write $f_{2}=e^{u+iv}$. Then the mapping $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$. Clearly $f$ is not spherically equivalent to a rational map if $f_{1}$ is not rational.
The map $f$ is at most points locally an embedding. We can compose with an automorphism on the source and target and assume that $f(0)=0$, $f^{\prime}(0)\not=0$, and that for some neighborhood $U\subset\mathbb{D}$ of 0 and a neighborhood $V\subset\mathbb{B}_{2}$ of 0, $f(U)\subset V$ is a properly embedded submanifold, $f|_{U}$ is one-to-one, and $f(\mathbb{D}\setminus U)\cap V=\varnothing$. Assume that $V$ is a ball. Consider a closed ball $B$ centered at the origin that is small enough so that $\overline{B}\subset V$ and so that $K=f^{-1}(B)$ is connected. The set $\mathbb{D}\setminus K$ is biholomorphic to some $\mathbb{A}_{1,r}$ for some $r$ via some $\varphi\colon\mathbb{A}_{1,r}\to\mathbb{D}\setminus K$. Then $f\circ\varphi$ is a proper map of $\mathbb{A}_{1,r}$ to $\mathbb{B}_{2}\setminus B=\mathbb{A}_{2,R}$ (where $R$ is the radius of $B$).
The set $\mathbb{D}\setminus K$ has real-analytic boundary and so $\varphi$ extends past the boundary. Therefore, $f\circ\varphi$ cannot extend past the boundary at any point on $\partial\mathbb{D}$ by the assumption on $f_{1}$ and so $f\circ\varphi$ is not rational.
Let us compare the effect of the automorphism groups for proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ versus $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form (9) below. For any affine subspace $L$ of dimension $m<n$ such that $L\cap\mathbb{B}_{n}\not$$\varnothing$, there exists a $\varphi\in\operatorname{Aut}(\mathbb{B}_{n})$ so that $\varphi(L)$ is given by $z_{m+1}=\cdots=z_{n}=0$.
The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group. In particular, given an affine subspace $L$ of dimension $m<n$ that intersects both the outer ball $\mathbb{B}_{n}$ and the inner ball $r\mathbb{B}_{n}$, elementary geometry says that there is a $\varphi\in\operatorname{Aut}(\mathbb{A}_{n,r})=U(n)$ so that $\varphi(L)$ is given by $z_{m+1}=c$ and $z_{m+2}=\cdots=z_{n}=0$ where $0\leq c<r$.
The effect of this difference on classification up to equivalence with respect to the embedding dimension $N_{f}$ is the following. For a proper ball map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ where $N_{f}<N$, there exists an automorphism $\psi\in\operatorname{Aut}(\mathbb{B}_{N})$ so that $\psi\circ f=F\oplus 0$ where $F\colon\mathbb{B}_{n}\to\mathbb{B}_{N_{f}}$ is a proper map ($N_{f}=N_{F}$). The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space, it can only “rotate” it. The proof is an elementary computation.
Proposition 2.4.
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N_{f}<N$. Then there exists a unitary $U\in U(N)$ and a unique $c\in[0,R)$ such that
(8) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0$
where $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{f},\sqrt{\frac{R^{2}-c^{2}}{1-c^{2}}}}$ is a proper map and $N_{f}=N_{F}$.
In the proposition, the 0 in the direct sum has $N-N_{f}-1$ components (so it can be left out if $N=N_{f}+1$). The proposition leads to classification of first degree proper rational maps. We first start with the equidimensional case. In this case, unitaries (so unitarily equivalent to the identity) are the only maps.
Proposition 2.5.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ is a proper rational map of degree 1 and $n\geq 2$, then $r=R$ and $f$ is a unitary map.
Proof.
We have that $f$ is an LFT that must take the unit sphere to the unit sphere, hence it is well-known that it must have the form
(9) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z,$
where $U$ is a unitary and $a\in\mathbb{B}_{n}$. We need to show that $a=0$. By precomposing with a unitary, we can assume that $a=(a_{1},0,\ldots,0)$ and
(10) $\phi_{a}(z)=\mathopen{}\mathclose{{\left(\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}},\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{2}}{1-z_{1}\bar{a}_{1}},\ldots,\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{n}}{1-z_{1}\bar{a}_{1}}}}\right).$
We note that the complex line $z_{2}=\cdots=z_{n}=0$ gets taken by $\phi_{a}$ to this same complex line, and on this line $\phi_{a}$ is simply the corresponding LFT in one dimension $\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$.
It remains to show that if $z_{1}\mapsto\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$ takes a circle of radius $r<1$ centered at zero to a circle of radius $R<1$ centered at zero. In other words, we must have that for all $\theta$, $\mathopen{}\mathclose{{\left\lvert a_{1}-re^{i\theta}}}\right\rvert^{2}=R^{2}\mathopen{}\mathclose{{\left\lvert 1-re^{i\theta}\bar{a}_{1}}}\right\rvert^{2}$. By expanding and equating coefficients one obtains that $a_{1}=0$ and $r=R$. In other words, $a=0$ and $f$ is a unitary. ∎
As a quick corollary we can remove the hypothesis of rationality and degree. By Hartogs phenomenon and the maximum principle, we find that a holomorphic proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself. That this map is a rational map of degree 1 then follows by Alexander’s theorem.
Corollary 2.6.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ is a proper holomorphic map and $n\geq 2$, then $r=R$ and $f$ is a unitary map.
When $N>n$, the result is a little more complicated.
Proposition 2.7.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1 and $N>n\geq 2$, then $R\geq r$ and $f$ is unitarily equivalent to the affine embedding
(11) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0.$
In the proof we will use the notion of a rank of a real polynomial $r(z,\bar{z})$ in $\mathbb{C}^{n}$. By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
(12) $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
where $\epsilon_{j}\pm 1$ and $P_{j}$ are holomorphic polynomials.
Proof.
Write $f=\frac{p}{q}$. As $f$ takes the unit sphere to the unit sphere we have
(13) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert q(z)}}\right\rVert^{2}=\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$
where $\lambda$ has to be a constant as $p$ and $q$ are of degree 1. Therefore, the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ has to be $n$ as that is the number of positive eigenvalues of the matrix of coefficients $\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$. In other words, $N_{f}=n$ and we can apply Proposition˜2.4. The $F$ in the proposition is unitarily equivalent to the identity by Proposition˜2.5. The values of the factors follow by elementary computation. ∎
We derive a few of basic results about the $k_{f}$ integer. First we note that for any rational map, it is enough to check an open dense subset of hyperplanes, that is, the general hyperplane. Similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication. The space of affine hyperplanes in $\mathbb{C}^{n}$ has the topology of the Grassmannian (minus the hyperplane at infinity).
Proposition 2.8.
If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map, then $\dim\operatorname{aff}\bigl(f(H)\bigr)=k_{f}$ for an open dense set of affine hyperplanes $H$.
The proof is to simply note that the affine dimension of $f(H)$ must be lower semicontinuous on the space of hyperplanes $H$, since the set of hyperplanes where the affine dimension is at most some dimension is defined by polynomial equations.
Next, we recall the proof of the classic result that $k_{f}\leq N-1$ for rational maps of spheres, as the technique will be useful later.
Proposition 2.9.
If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
Proof.
Let $\langle\cdot,\cdot\rangle$ denote the standard sesquilinear inner product so that the sphere $S^{2n-1}$ is defined by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}=\langle z,z\rangle=1$. For any fixed $w\in\mathbb{C}^{n}$, define the Segre variety
(14) $\Sigma_{w}=\{z\in\mathbb{C}^{n}:\langle z,w\rangle=1\}.$
Similarly if $\omega\in\mathbb{C}^{N}$, we have the Segre variety of the target sphere $\Sigma_{\omega}\subset\mathbb{C}^{N}$. Note that the Segre variety for a sphere is an affine hyperplane. If $f=\frac{p}{q}$, then the fact that $f$ takes the sphere to the sphere means that $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}=g(z,\bar{z})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1}}\right)$, or $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}-1=\frac{g(z,\bar{z})}{\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1}}\right)$, where $g$ is a polynomial. By polarizing this identity, we find the equality
(15) $\mathopen{}\mathclose{{\left\langle f(z),f(w)}}\right\rangle-1=\frac{g(z,\bar{w})}{q(z)\overline{q(w)}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right).$
Writing $P=q^{-1}(0)$ for the pole set of $f$, we therefore have that $f(\Sigma_{w}\setminus P)\subset\Sigma_{f(w)}$ for all $w$ where $f(w)$ is defined. As any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$. Therefore, there is an open set of affine hyperplanes which $f$ takes into hyperplanes. Therefore $k_{f}\leq N-1$. ∎
The $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
Proposition 2.10.
For a map $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$, $\phi$ an LFT on $\mathbb{C}^{n}$, and $\psi$ an LFT on $\mathbb{C}^{N}$, we have $k_{f}=k_{\psi\circ f\circ\phi}$. In other words, $k_{f}$ is both a spherical and a unitary invariant.
As part of Theorem˜1.2, we claim that $k_{H_{d}}=N-1$. Let us prove that now.
Proposition 2.11.
For the homogeneous map $H_{d}$, we have $k_{H_{d}}=N-1$ where $N=\binom{n+d-1}{d}$ is the target dimension of $H_{d}$.
Proof.
The map $H_{d}$ has as its components every degree $d$ monomial in $n$ variables, $H_{d}=(z_{1}^{d},\sqrt{d}\,z_{1}^{d-1}z_{2},\ldots)$. If we consider what $H_{d}$ does to the hyperplane $\{z_{1}=c\}$ for nonzero constants $c$, we find that except for the first component which becomes the constant $c^{d}$, all the other components are affine independent. In other words, the affine span of $H_{d}(\{z_{1}=c\})$ is of dimension N1N-1. ∎
Next we consider the $k_{f}$ integer for juxtapositions. We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [dangelo-2009-complexity]. The next proposition therefore shows that for rational ball (or annulus) maps, $k_{f}=N_{f}-1$ happens only on the non-extremal points of this convex set.
Proposition 2.12.
Suppose that $F_{t}$ is a juxtaposition of two rational proper maps of balls (or annuli) $F_{0}$ and $F_{1}$, that is, $F_{t}=\sqrt{1-t}\,F_{0}\oplus\sqrt{t}\,F_{1}$ for some $t\in(0,1)$, and further suppose that $\mathopen{}\mathclose{{\left\lVert F_{0}(z)}}\right\rVert^{2}\not\equiv\mathopen{}\mathclose{{\left\lVert F_{1}(z)}}\right\rVert^{2}$. Then $k_{f}\leq N_{f}-2$.
The condition on the norms is to ensure that we are taking the convex combination of a point with itself in the space of squared norms. The condition is equivalent to requiring there is no unitary $U$ such that $F_{1}=UF_{0}$ (possibly adding zero components to make the target dimensions equal).
Proof.
We note that if $k_{F_{0}}\leq N_{F_{0}}-2$ or $k_{F_{1}}\leq N_{F_{1}}-2$, then we are done trivially. Therefore suppose that $k_{F_{0}}=N_{F_{0}}-1$ and $k_{F_{1}}=N_{F_{1}}-1$. Also note from the proof of Proposition˜2.9 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies. In particular, we find these vectors in an $N_{f}$-dimensional hyperplane which we can assume is the entire target space. Apart from these facts that we can simply treat these as rational maps. The integers $k_{f}$ and $N_{f}$ and the fact about the normal vectors are all invariant under any LFT and so we only need to prove the result for $F_{0}\oplus F_{1}$. We can also assume that $F_{0}$ has $N_{F_{0}}$ components and $F_{1}$ has $N_{F_{1}}$ components. We can further homogenize both maps up to the same degree and instead of affine hyperplanes we deal with linear subspaces. The dimensions increase by 1, but that is not relevant to us as we are interested in their difference. In other words, what we have is two homogeneous polynomial maps of the same degree $F_{0}$ and $F_{1}$, we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent. We further have that for any codimension 1 subspace $H$, the span of $F_{0}(H)$ is of codimension 1 (in the codomain of $F_{0}$) and same for $F_{1}(H)$. It is possible that components of $F_{0}$ and $F_{1}$ are not linearly independent when put together. After composing with a linear we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components. Write the target space as $\mathbb{C}^{j}\times\mathbb{C}^{k}\times\mathbb{C}^{\ell}$. The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ ($v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, $0\in\mathbb{C}^{\ell}$) so that $\xi\cdot v=0$ for all $\xi\in P(H)$. There is also a vector $w=(w_{1},0,w_{3})$ so that $\xi\cdot w=0$ for all $\xi\in P(H)$. We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, we found that $P(H)$ lives in a codimension 2 space and we are done. Thus suppose that there exists a $v=w=(v_{1},0,0)$ for every subspace $H$. But our assumption about the normal vectors of the hyperplanes would give a contradiction. Thus $v$ and $w$ are always linearly independent and we find that $N_{P}-k_{p}\geq 2$. ∎
We have seen that the homogeneous map $H_{d}$ is an annulus map. We also have that juxtapositions of homogeneous maps and their affine embeddings are annulus maps. Here is an example that is not constructed in this way from homogeneous maps.
Example 2.
With the construction introduced in [helal-2025-proper], one finds the following map
(16) $f(z_{1},z_{2})=\mathopen{}\mathclose{{\left(\tfrac{2}{\sqrt{5}}z_{1}^{3},\tfrac{2\sqrt{2}}{\sqrt{5}}z_{1}^{2}z_{2},\tfrac{2}{\sqrt{5}}z_{1}z_{2}^{2},z_{1}z_{2},z_{2}^{2},\tfrac{1}{\sqrt{5}}z_{1}}}\right).$
This map gives a proper map of $\mathbb{A}_{2,\frac{1}{2}}$ to $\mathbb{A}_{6,\frac{1}{4}}$. It takes the unit sphere to the unit sphere and the half sphere to quarter sphere. It takes no other sphere centered at the origin to a sphere centered at the origin. It is not unitarily equivalent to a juxtaposition of several homogeneous maps, nor is it an affine embedding of such a map. If it were, then it would take all spheres centered at the origin to spheres centered at the origin. We have that $N_{f}=6$ and one can compute that the $k_{f}$ number is $3$, so $k_{f}<N_{f}-1$. To see why it is $3$, note that if we restrict to a 1-dimensional affine space $H$, the map becomes a one variable cubic polynomial and so the affine span can be at most $3$. That it is exactly $3$ can be seen by restricting to a hyperplane making $z_{2}$ constant.
We remark that while Theorem˜1.2 implies that for annuli maps, $k_{f}=N-1$ means $f$ is homogeneous, there must be extremal points with $k_{f}<N-1$ in the convex set describing the rational proper annulus maps as the example above shows that there exist maps that are not constructed out of homogeneous maps via juxtapositions.
The construction from [helal-2025-proper] can be used to obtain cubic examples in higher dimensions and also quadratic rational maps with the same properties (not taking more than two zero-centric spheres to zero-centeric spheres, so not constructed out of homogeneous maps).
3. First gap
Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [dangelo-2012-pfisters]*Theorem 1.1. We only need the quadratic version of this result, but we state the more general result. Recall that the rank of a real polynomial $r(z,\bar{z})$ is the rank of its matrix of coefficients, or alternatively the smallest $k$ such that $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2}$.
Theorem 3.1.
Let $n\geq 1$, $s_{1},s_{2}$ nonzero real numbers of the same sign, and $Q(z,\bar{z})$ be a nonzero real polynomial on $\mathbb{C}^{n}$. Then
(17) $\operatorname{rank}Q(z,\bar{z})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})\geq\operatorname{rank}(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})=\binom{n+2}{2}$.
The proof follows by roughly the same argument as in [dangelo-2012-pfisters] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors. What must be newly proved here is a “diagonal version” of the theorem. We only have this proof for two factors, which is sufficient for our purposes. We conjecture that the result holds for any number of factors. Note that the result is not true as stated if the numbers $s_{j}$ have opposite signs. For example, $(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}+1)(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}-1)=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{4}-1$ is of rank 2 and not $\binom{n+2}{2}=\binom{3}{2}=3$.
Proof.
The main new claim that must be proved is the “diagonal version” of this theorem. That is, for a real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
(18) $\bigl(Q(x)(x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)\\ \geq\#\bigl((x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)=\binom{n+2}{2}$,
where #\# means the number of nonzero coefficients in the given polynomials when they are expanded. We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{d}$ and applying the claim for positive $s_{1},s_{2}$.
First suppose that $n=1$. For this case, assume $s_{1}$ and $s_{2}$ are both negative. The polynomial
(19) $R(x)=Q(x)(x+s_{1})(x+s_{2})$
has at least $2$ positive real roots. By Descartes’ rule of signs, there must be at least $2$ sign changes among the coefficients of the polynomial and therefore there must be at least $3$ nonzero coefficients, which is precisely the claimed result for $n=1$.
Next, let us suppose $n=2$. Again let $R(x)=Q(x)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$. Write $R(x)=x_{1}^{j}x_{2}^{k}\widetilde{R}(x)$ (and similarly $Q(x)=x_{1}^{j}x_{2}^{k}\widetilde{Q}(x)$) where $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ are both not identically zero. Applying, the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively). One of these terms can be the constant, and therefore, we have $1+2n$ terms in $\widetilde{R}(x)$ which are all pure powers. Next we will show that there is a term in $\widetilde{R}(x)$ that involves both $x_{1}$ and $x_{2}$. Note that the vector field $L=\frac{\partial}{\partial x_{1}}-\frac{\partial}{\partial x_{2}}$ is tangent to each $(x_{1}+x_{2}+s_{j})$. Therefore, if we apply $L$ $\nu$ number of times, we get
(20) $L^{\nu}\widetilde{R}(x)=\bigl(L^{\nu}\widetilde{Q}(x)\bigr)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2}).$
As $\widetilde{Q}$ is a polynomial, $L^{\nu}\widetilde{Q}$ must be identically zero for some $\nu$, so take $\nu$ such that $L^{\nu-1}Q$ is not zero but $L^{\nu}Q\equiv 0$. Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation and so it is a function $x_{1}+x_{2}$, that is $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$ $S$ is not identically zero. But then
(21) $L^{\nu-1}\widetilde{R}(x)=S(x_{1}+x_{2})(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$
is a polynomial in $(x_{1}+x_{2})$ of degree at least 2 and hence it must have a mixed monomial. Since $L^{\nu-1}\widetilde{R}(x)$ has a mixed monomial, then so does $\widetilde{R}(x)$ as $L$ cannot manufacture mixed monomials out of pure ones. Therefore, $R(x)$ has at least $1+2\cdot 2+1=\binom{2+2}{2}=6$ terms.
Moreover, we guarantee that there are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$, and at least 3 terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m\geq k$. Note that also there are no terms at all with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$. Similarly, there are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$, and at least 3 terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell>k$ and $m\geq j$. And no terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell<k$.
We now attack the general problem for $n>2$. Pick any pair of variables, without loss of generality it is $x_{1}$ and $x_{2}$. Write $x=(x_{1},x_{2},x^{\prime})$ and fix one of the standard monomial orders on the remaining variables $x^{\prime}$, such that multiplying a monomial in $x^{\prime}$ by any one of the variables $x_{3},\ldots,x_{n}$ moves the monomial further in the ordering. Using multinomial notation, write
(22) $R(x)=\sum_{\alpha}r_{\alpha}(x_{1},x_{2})(x^{\prime})^{\alpha}\quad\text{and}\quad Q(x)=\sum_{\alpha}q_{\alpha}(x_{1},x_{2})(x^{\prime})^{\alpha}.$
we find the smallest monomial $(x^{\prime})^{\mu}$ in the given monomial order for which the $q_{\alpha}$ is not identically zero. We then have that
(23) $r_{\mu}(x_{1},x_{2})=q_{\mu}(x_{1},x_{2})(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2}).$
All the monomials of $r_{\mu}(x_{1},x_{2})(x^{\prime})^{\mu}$ must appear in $R(x)$ by minimality of $\mu$. We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these $a_{\ell m}$ is as in the “Moreover” remark above. We use this this remark for every pair of variables to count monomials. For every pair there is one monomial that we might be also counting for other pairs, so we count it only once. Then there are two monomials corresponding to each variable that may appear in the count for every pair that includes this variable. Finally, for every pair there is at least one monomial that does not appear in the count for every other pair. In other words, We have at least
(24) $1+2n+\binom{n}{2}=\binom{n+2}{n}$
distinct monomials in $R(x)$. The claim is proved.
Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [dangelo-2012-pfisters]. That is, there exists some ordering of monomials so that if we write out the matrix of coefficients of the expressions involved, multiplication by $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ moves along diagonals. The rank of the matrix of coefficients is the rank of the function. The rank of a matrix can be estimated from below by considering the number of nonzero terms in the superdiagonal that is furthest up and to the right that has nonzero terms. Up to multiplication with a monomial, the terms on this diagonal are polynomials in $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$, and so by replacing $x_{j}=\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ we obtain the situation of the claim and the proof follows. See [dangelo-2012-pfisters] for more details. ∎
We can now prove Theorem˜1.1, that is, if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ ($n\geq 2$) is a proper holomorphic map and $n<N<\binom{n+1}{2}$, then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$.
Proof of Theorem˜1.1.
By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$. Lemma 3.11 from [helal-2025-proper] gives us
(25) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\Bigr)\cdot\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q_{2}(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
for a real polynomial $Q_{2}(z,\bar{z})$.
If $Q_{2}\equiv 0$, then by the argument in the proof of Theorem 3.12 from [helal-2025-proper], $f$ is a polynomial, that is, $q\equiv 1$. Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$). Therefore, $\deg f=1$. Proposition˜2.7 implies that $f$ is an affine embedding and that $R\geq r$.
Suppose $Q_{2}\not\equiv 0$, we will get a contradiction. By applying Proposition˜2.4, we reduce to the case when $N=N_{f}$. The dimension $N$ is then the number of components of $f$, which is equal to the number of components of $p$, which we can assume are linearly independent as $N_{f}=N$. Therefore, $N$ is equal to the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$. Theorem˜3.1 gives us
[…]
In the inequalities we have used the facts that $\operatorname{rank}(A+B)\geq-(\operatorname{rank}A)+\operatorname{rank}B$ and that $\operatorname{rank}\bigl(\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}A(z,\bar{z})\bigr)=\operatorname{rank}A(z,\bar{z})$, both of which follow immediately from the definition of rank. By hypothesis, we find a contradiction as we assumed $N_{f}<\binom{n+1}{2}$.
We therefore obtained $N_{f}=n$ and that the degree is $1$. The conclusion follows from Proposition˜2.5 or Proposition˜2.7 depending on whether $N>n$ or $N=n$. The interval is sharp; when $N=\binom{n+1}{2}$, then besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding. ∎
We should remark that for specific $\mathbb{A}_{n,r}$ and $\mathbb{A}_{\binom{n+1}{2},R}$ we either get the affine embeddings if $R\geq r$, or we get the homogeneous map $H_{2}$ if $R=r^{2}<r$. For general $n$, it is not clear if other maps exist when $N=\binom{n+1}{2}$, for perhaps different $R$. For $n=2$, the homogeneous $H_{2}$ and the affine embeddings are the only proper maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{\binom{2+1}{2},R}=\mathbb{A}_{3,R}$. See Theorem˜1.4.
4. Homogeneous maps via the generic hyperplane rank
To prove the theorem on classification of homogeneous maps, we will prove a more general result about rational maps taking two spheres to spheres. To state the result, it may be easiest to work in the projective space $\mathbb{P}^{n}$. The automorphism group $\operatorname{Aut}(\mathbb{P}^{n})$ of $\mathbb{P}^{n}$ is the set of linear fractional transformations or LFTs. Consider some embedding $\mathbb{C}^{n}\subset\mathbb{P}^{n}$ and then we can consider the sphere $S^{2n-1}$ as a subset of $\mathbb{P}^{n}$. When we refer to affine coordinates, we refer to this embedding and the coordinates $(z_{1},\ldots,z_{n})$. We say $\mathcal{S}$ is an LFT sphere if there exists an LFT $\varphi\in\operatorname{Aut}(\mathbb{P}^{n})$ so that $\mathcal{S}=\varphi(S^{2n-1})$.
Our first result is about rational maps taking two LFT spheres to LFT spheres. The definition of $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$. In the proof we will work in homogeneous coordinates on the projective space, and so a general hyperplane in $H\subset\mathbb{P}^{n}$ is an $n$-dimensional linear subspace in $\mathbb{C}^{n+1}$.
We use $z=(z_{0},\ldots,z_{n})$ for the homogeneous coordinates on $\mathbb{P}^{n}$. To write a rational function $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
The sphere $S^{2n-1}\subset\mathbb{C}^{n}\subset\mathbb{P}^{n}$ in our chosen affine embedding when put into homogeneous coordinates is given by the equation
(27) $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0,\quad\text{where}\quad J_{n}=\operatorname{diag}(-1,1,\ldots,1)=\begin{bmatrix}-1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
[…]
(28) $\mathopen{}\mathclose{{\left\langle T^{*}J_{n}Tz,z}}\right\rangle=0,$
for an invertible $T$, that is, it is $\mathopen{}\mathclose{{\left\langle Az,z}}\right\rangle=0$ for some hermitian $(n+1)\times(n+1)$ matrix $A$ with $n$ positive and $1$ negative eigenvalues. If $\varphi$ is the LFT denoted by the matrix $T$, then $\mathopen{}\mathclose{{\left\langle T^{*}J_{n}Tz,z}}\right\rangle=0$ represents the sphere $\varphi^{-1}(S^{2n-1})$. We can also do the same in $\mathbb{C}^{N+1}$ on the target, using instead the matrix $J_{N}$.
Theorem 4.1.
Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to two LFT spheres and $k_{f}=N-1$. Then $f\circ\alpha=\beta\circ f$ for some LFTs $\alpha\in\operatorname{Aut}(\mathbb{P}^{n})$ and $\beta\in\operatorname{Aut}(\mathbb{P}^{N})$ neither of which is the identity.
More precisely, using the coordinates above, if one of the spheres on the source (resp. the target) are given by $J_{n}$ (resp. $J_{N}$) and the other by a hermitian matrix $A$ (resp. $B$), then $\alpha$ is represented by the matrix $J_{n}A$ and $\beta$ is represented by the matrix $J_{n}B$.
The main idea behind the proof is that $k_{f}=N-1$ means that the map $f$ induces a map of the corresponding Grassmannians. To figure out the map one uses the Segre variety of the LFT spheres. But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different mapping of Grassmannians. Comparing the two maps gives us the corresponding $\alpha$ and $\beta$.
Proof.
As we said we will work in homogeneous coordinates, that is, in $\mathbb{C}^{n+1}$ and $\mathbb{C}^{N+1}$. Starting with a rational map $f$, we clear denominators and homogenize to represent $f$ by a homogeneous polynomial map $F=(F_{0},F_{1},\ldots,F_{N})\colon\mathbb{C}^{n+1}\to\mathbb{C}^{N+1}$ where $F_{0}$ is the denominator. After applying an LFT we can assume that one of the matrices on both the source and the target is the standard sphere $S^{2n-1}$ and $S^{2N-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$. For any hermitian matrix $M$ and a point $w$, we write the Segre variety of the homogenized LFT sphere given by $\mathopen{}\mathclose{{\left\langle Mz,z}}\right\rangle=0$ as
(29) $\Sigma_{w}^{M}=\{z:\mathopen{}\mathclose{{\left\langle Mz,w}}\right\rangle=0\}.$
In affine coordinates, the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z),p(w)}}\right\rangle-q(z)\overline{q(w)}=g(z,\bar{w})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right)$ for a polynomial $g$. In homogeneous coordinates, this translates to
(30) $\mathopen{}\mathclose{{\left\langle J_{N}F(z),F(w)}}\right\rangle=G(z,\bar{w})\mathopen{}\mathclose{{\left\langle J_{n}z,w}}\right\rangle.$
This equation implies that
(31) $F(\Sigma_{w}^{J_{n}})\subset\Sigma_{F(w)}^{J_{N}}.$
We are assuming that we have a homogenized LFT sphere in $\mathbb{C}^{n+1}$ represented by the $(n+1)\times(n+1)$ hermitian matrix $A$ that $F$ takes to a homogenized LFT sphere in $\mathbb{C}^{N+1}$ represented by the $(N+1)\times(N+1)$ hermitian matrix $B$. Therefore, we also have
(32) $F(\Sigma_{w}^{A})\subset\Sigma_{F(w)}^{B}.$
Let $S=(AJ_{n}^{-1})^{*}=J_{n}A$. As
(33) $\mathopen{}\mathclose{{\left\langle Az,w}}\right\rangle=\mathopen{}\mathclose{{\left\langle AJ_{n}^{-1}J_{n}z,w}}\right\rangle=\mathopen{}\mathclose{{\left\langle J_{n}z,(AJ_{n}^{-1})^{*}w}}\right\rangle,$
we find that $\Sigma_{w}^{A}=\Sigma_{Sw}^{J_{n}}$, and so
(34) $F(\Sigma_{w}^{A})\subset\Sigma_{F(w)}^{B}\quad\text{and}\quad F(\Sigma_{w}^{A})=F(\Sigma_{Sw}^{J_{n}})\subset\Sigma_{F(Sw)}^{J_{N}}.$
The condition that $k_{f}=N-1$ means in homogeneous coordinates that given a linear $n$-dimensional subspace $V\subset\mathbb{C}^{n+1}$, the linear span of $F(V)$ in $\mathbb{C}^{N+1}$ is $N$-dimensional. The Segre variety $\Sigma_{w}^{A}$ is an $n$-dimensional subspace, and $\Sigma_{F(w)}^{B}$ and $\Sigma_{F(Sw)}^{J_{N}}$ are $N$-dimensional subspaces. In other words, we find that
(35) $\Sigma_{F(Sw)}^{J_{N}}=\Sigma_{F(w)}^{B}.$
That is, using coordinates $(Z_{0},\ldots,Z_{N})$ in $\mathbb{C}^{N+1}$, we have that $\mathopen{}\mathclose{{\left\langle J_{N}Z,F(Sw)}}\right\rangle=0$ and $\mathopen{}\mathclose{{\left\langle BZ,F(w)}}\right\rangle=0$ denote the same subspace. As this is true for every $w$, we have that $\mathopen{}\mathclose{{\left\langle J_{N}Z,F(Sw)}}\right\rangle=\lambda\mathopen{}\mathclose{{\left\langle BZ,F(w)}}\right\rangle$, where $\lambda$ must be a constant as the two functions are both degree-$1$ homogeneous in $Z$ and degree-$d$ homogeneous in $w$ (assuming $F$ is of degree $d$). We can rewrite this equation as $J_{N}^{*}F(Sw)=\lambda B^{*}F(w)$, or in other words,
(36) $F(Sw)=\lambda J_{N}^{-*}B^{*}F(w).$
Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$, we have
(37) $F(Sw)=\lambda TF(w).$
If we take $\alpha$ to be the LFT given by the matrix $S$ and $\beta$ be the LFT given by the matrix $T$, then we obtain in affine coordinates that
(38) $f\circ\alpha=\beta\circ f,$
as required. The formulas for $\alpha$ and $\beta$ follow. ∎
In particular, $f$ now also takes the LFT sphere $\alpha^{-1}(S^{2n-1})$ to the LFT sphere $\beta^{-1}(S^{2N-1})$. It turns out that these spheres need not be the ones we started with. By iterating, we find that for any $k\in\mathbb{Z}$ we have that
(39) $f\circ\alpha^{k}=\beta^{k}\circ f.$
So we also get that $f$ takes $\alpha^{-k}(S^{2n-1})$ to $\beta^{-k}(S^{2N-1})$ for any $k$. In our particular case, when all spheres involved are centered at zero, $f$ takes infinitely many spheres centered at zero of various radii to spheres centered at zero. In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, then we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
Lemma 4.2.
Suppose $n\geq 2$ and $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map such that $f(S^{2n-1})\subset S^{2N-1}$ and for some $r,R<1$ we have that $f(rS^{2n-1})\subset RS^{2N-1}$. Suppose also that $k_{f}=N-1$. Then there are two sequences of distinct numbers $\{r_{k}\}$ and $\{R_{k}\}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{K}S^{2N-1}$ for every $k$.
Proof.
We set things up as in the proof of Theorem˜4.1, and so the matrices $A$ and $B$ are given as
(40) $A=\begin{bmatrix}-r^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}\quad\text{and}\quad B=\begin{bmatrix}-R^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
Then
(41) $S=J_{n}A=\begin{bmatrix}r^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}\quad\text{and}\quad T=J_{N}B=\begin{bmatrix}R^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
As we said what we have is that $f$ takes $\alpha^{-1}(S^{2n-1})$ to $\beta^{-1}(S^{2N-1})$. In affine coordinates, $\alpha$ is simply a scaling by $\frac{1}{r^{2}}$ and $\beta$ is a scaling by $\frac{1}{R^{2}}$, and so $\alpha^{-1}(S^{2n-1})=r^{2}S^{2n-1}$ and $\beta^{-1}(S^{2N-1})=R^{2}S^{2N-1}$. By iterating we get the desired sequence. ∎
We are now ready to prove Theorem˜1.2 using the results of our previous work in [helal-2025-proper].
Proof of Theorem˜1.2.
One direction of the theorem is clear and follows from the discussion so far. What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$. By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜4.2. Therefore, $f$ takes infinitely many spheres centered at zero to spheres centered at zero. In [helal-2025-proper], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros. So for some unitary $U$, we can assume that
(42) $f=U(c_{1}H_{d_{1}}\oplus\cdots c_{\ell}H_{d_{\ell}}\oplus 0),$
where we can assume that $d_{1},\ldots,d_{\ell}$ are distinct. Since we are assuming that $k_{f}=N-1$, we must have that $N_{f}=N$ and hence there are no zero components. Then, via the argument in the proof of Proposition˜2.11, we can see that $k_{f}=N$-$\ell$. Hence $\ell=1$, and $f=UH_{d}$ as required. ∎
Let us now prove Theorem˜1.4 as an application of Theorem˜1.2. We already have that proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$ are rational. If the degree of the map is 1, then via Proposition˜2.7, it is unitarily equivalent to the affine embedding
(43) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{1},\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{2},\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}}}\right),$
where $R\geq r$. In this case, clearly $N_{f}=2$ and $k_{f}=1$.
The theorem will therefore follow if we can show that any map of degree greater than 1 is unitarily equivalent to $H_{2}$. This result follows by Faran’s [faran-1982-maps] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The following lemma finishes the proof of Theorem˜1.4.
Lemma 4.3.
Suppose that $f\colon\mathbb{A}_{2,r}\to\mathbb{A}_{3,R}$ is a rational proper map of degree greater than 1. Then $f$ is unitarily equivalent to the homogeneous map $H_{2}$.
Proof.
The map $f$ extends to a proper map of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. Faran’s result implies that a proper rational map of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ is spherically equivalent to one of four maps: the linear embedding $(z_{1},z_{2},0)$, the Whitney map $(z_{1},z_{1}z_{2},z_{2}^{2})$, the homogeneous $(z_{1}^{2},\sqrt{2}\,z_{1}z_{2},z_{2}^{2})$, and the map $(z_{1}^{3},\sqrt{3}\,z_{1}z_{2},z_{2}^{3})$. The issue is that the automorphisms used in Faran’s classification are automorphisms of the ball and these do not preserve the inner sphere boundary of the annulus. Spherical equivalence does, however, preserve the degree of the map and the integer $k_{f}$. For the three maps of degree greater than 1, we find $k_{f}=2=N-1$. To compute $k_{f}$, it is sufficient to consider the hyperplanes $z_{2}=\text{constant}$. In other words, if we have an annulus map as in the statement and of degree greater than 1, then $k_{f}=2=N-1$. Hence Theorem˜1.2 says that $f$ is unitarily equivalent to $H_{2}$. ∎
5. Other remarks on proper maps of annuli
Let us put together some further minor remarks. Firstly, there are two other domains in $\mathbb{C}^{n}$ whose automorphism group is $U(n)$. There is the punctured ball $\mathbb{B}_{n}\setminus\{0\}$ and the complement of the closed unit ball $\mathbb{C}^{n}\setminus\overline{\mathbb{B}_{n}}$. For maps from a punctured ball to a punctured ball, the maps must extend through the origin by the Riemann extension theorem if $n=1$ or via the Hartogs phenomenon if $n>1$. We are then simply dealing with proper maps of balls that take the origin to the origin. Since the automorphism group of the ball is transitive, every spherical equivalence class of proper maps of balls contains a map taking 0 to 0. The case of the complement of the closed unit ball was studied in [helal-2025-proper]. In that case, for $n\geq 2$, the proper maps are polynomial sphere maps whose norm goes to infinity at infinity. In either case, the classification is different in these degenerate cases, and in both cases there are more maps than for annuli.
Another remark we wish to make is that we have already seen that in the first interval we have the restriction $r\leq R<1$ for a map to exist. A question may arise about which $R$ admit a proper map in general. It is relatively simple to see that as long as we do not place any restriction on $N$ we may get any $r$ and $R$ whatsoever. That is, given $n$, $r$, and $R$, pick $d$ so that $r^{d}\leq R$. Then as $H_{d}$ takes the $r$-sphere to the $r^{d}$-sphere we have that the map
(44) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2d}}}\,H_{d}(z)\oplus\sqrt{\frac{R^{2}-r^{2d}}{1-r^{2d}}}$
is a proper map of $\mathbb{A}_{n,r}$ to $\mathbb{A}_{N,R}$ where $N=\binom{n+d-1}{d}+1$.
The previous discussion suggests that the existence of a rational proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ of degree $d$ may imply a relationship between $r$, $R$, and $d$. For homogeneous maps as above we have $r^{d}\leq R$, with $H_{d}$ itself achieving the equality in this bound. Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general. If one also requires that $N$ is the embedding dimension ($N=N_{f}$), then more can be said. In this case, if $d=1$, then $R=r$. For higher degree, one could use juxtaposition to get a range. For example, if we juxtapose $H_{1}$ and $H_{d}$, then via the calculation in the introduction, we have that for $t\in(0,1)$, the map $\sqrt{1-t}\,H_{1}\oplus\sqrt{t}\,H_{d}$ has $R=\sqrt{(1-t)r^{2}+tr^{2d}}$, and so any $R$ in the range between $r^{d}$ and $r$ is achieved by these maps. So the question could be modified for the case when $N=N_{f}$ and degree $d>1$, to ask if $R$ equal to or bigger than $r$ is possible.
Finally, we remark that given our results and our examples, it is plausible to conjecture a degree bound in terms of the dimensions. Some bound must exist by the theorem of Forstnerič, as annulus maps extend to proper rational maps of balls. However, in the annulus case, it appears that the homogeneous map achieves the highest degree for a given set of dimensions. That is, we conjecture that as long as $n\geq 2$, then $N\geq\binom{n+d-1}{d}$; note that this is a bound for $d$, but it does not have a nice formula. All the other maps we know how to construct are either juxtapositions of homogeneous maps or have lower degree than the homogeneous for a given target dimension.

### v2

@@
 Abstract.
-We study rational proper holomorphic maps of annuli in complex euclidean spaces, that is, domains with $U(n)$ as the automorphism group.
+We study proper holomorphic maps of annuli in complex Euclidean spaces, that is, domains with $U(n)$ as the automorphism group.
 By the Hartogs phenomenon and a result of Forstnerič, such maps are always rational and extend to proper maps of balls.
@@
 Next we find a necessary and sufficient condition for a map to be homogeneous: A proper map of annuli is homogeneous if and only if its general hyperplane rank, the affine dimension of the image of a general hyperplane, is exactly N1N-1.
-This result is really a result classifying homogeneous proper maps of balls.
-A homogeneous proper ball map takes spheres centered at the origin to spheres centered at the origin.
+As a corollary, we obtain a classification of homogeneous proper maps of balls.
+A homogeneous proper ball map takes all spheres centered at the origin to spheres centered at the origin.
 We show that if a proper ball map has general hyperplane rank N1N-1 and takes one sphere centered at the origin to a sphere centered at the origin, then it is homogeneous.
 Another corollary of this result is a complete classification of proper maps of annuli from dimension 2 to dimension 3.
+Finally, we give a complete normal form of rational proper maps of annuli of degree 2.
 Key words and phrases:
@@
 1. Introduction
-Let $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ denote the unit ball.
-A well-studied problem in complex analysis is to attempt to classify the proper holomorphic maps $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$.
-If $n=N=1$, then by a classical theorem of Fatou [fatou-1923-fonctions], $f$ is a finite Blaschke product and hence rational.
-If $n=N>1$, then by a theorem of Alexander [alexander-1977-proper], $f$ is an automorphism of the unit ball.
-No maps exist if $N<n$ for elementary reasons.
-Dor and Løw [dor-1990-proper, low-1985-embeddings] show that if $N>n>1$, there exist non-rational maps, even in the case of codimension 1, that is, $N=n+1$.
-If in the case $n>1$ one assumes some regularity on the boundary, then by a theorem of Forstnerič [forstneric-1989-extending], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$.
-Faran [faran-1982-maps] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are equivalent to one of four polynomial maps.
-For maps that are $C^{3}$ up to the boundary, Webster [webster-1979-mapping] showed that for $n\geq 3$, there is no map from $\mathbb{B}_{n}$ to $\mathbb{B}_{n+1}$ that is not equivalent up to automorphisms (spherically equivalent) to the linear embedding $z\oplus 0$.
-Faran [faran-1986-linearity] proved the same conclusion (only the linear embedding) for maps real-analytic up to the boundary for all $N$ up to N=2n2N=2n-2.
-Huang [huang-1999-linearity] improved the regularity to $C^{2}$.
-For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding.
-The map $z^{\otimes d}$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$.
-This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary mapping to what we will call $H_{d}$ direct sum a number of zero components.
-Rudin [rudin-1984-homogeneous] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$.
-See D’Angelo [dangelo-1988-polynomial] for a simpler proof of this fact.
-In general, see the books by D’Angelo [dangelo-1993-several, dangelo-2019-hermitian, dangelo-2021-rational] for a more thorough treatment of proper maps of balls.
-The ball can be regarded as the strongly pseudoconvex domain with the most symmetry.
-Let us consider then a non-pseudoconvex domain with a lot of symmetry, the annulus.
+One of the fundamental questions in complex analysis is to understand proper holomorphic maps $f\colon U\to V$ between two domains $U\subset\mathbb{C}^{n}$ and $V\subset\mathbb{C}^{N}$.
+In general, few if any such maps exist unless $U$ and $V$ are related.
+If $U$ and $V$ are the same type of domains and have many symmetries, then many such maps exist in general.
+The domain with the “largest” automorphism group is the unit ball $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ and the space of proper maps of balls $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ has a rich structure and has been extensively studied.
+We will focus on proper maps of annuli, as annuli are domains with a large automorphism group.
 Write $\mathbb{A}_{n,r}=\mathbb{B}_{n}\setminus\overline{r\mathbb{B}_{n}}=\{z\in\mathbb{C}^{n}:r<\mathopen{}\mathclose{{\left\lVert z}}\right\rVert<1\}$ for an annulus, where we will always assume that $0<r<1$.
@@
 When $n=1$ there are also the automorphisms $e^{i\theta}\frac{r}{z}$.
-The same classical argument to classify the automorphism group $\mathbb{A}_{1,r}$ can be extended to show that in one dimension a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case it is either equal to $e^{i\theta}z^{d}$ or $r^{d}e^{i\theta}z^{-d}$.
-In dimension $n\geq 2$, due to the inner pseudoconcave boundary, via the Hartogs phenomenon and the theorem of Forstnerič on rationality of proper ball maps, all proper holomorphic maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are rational, see Theorem˜2.2, but they can be more complicated.
-In fact, holomorphic proper maps of annuli are in one to one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero.
-See [helal-2025-proper] for more details.
-For $z\in\mathbb{C}^{n}$, if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
-(1) $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
+Notice that unlike the ball, the annulus is non-pseudoconvex; its boundary consists of a strongly pseudoconvex part and a strongly pseudoconcave part.
+It is not difficult to show that if $n>1$, then proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ extend to proper maps of balls that also take the $r$-sphere to the $R$-sphere.
+Their study therefore involves proper maps of balls.
+Our goal is to prove results for annuli analogous to the known results for balls and to contrast where these results fundamentally differ.
+We therefore very briefly discuss some of the known results for balls.
+Consider then a proper holomorphic map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$.
+If $n=N=1$, then by a classical theorem of Fatou [16], $f$ is a finite Blaschke product and hence rational.
+If $n=N>1$, then by a theorem of Alexander [1], $f$ is an automorphism of the unit ball.
+Dor [13] and Løw [22] show that if $N>n>1$, there exist nonrational maps, even in the case of codimension 1, that is, $N=n+1$.
+If $n>1$ and the map is sufficiently smooth up to the boundary, then by a theorem of Forstnerič [17], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$.
+Faran [14] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are spherically equivalent (up to composition with automorphisms of the balls) to one of four polynomial maps.
+For $n\geq 3$, by a series of papers of Webster [24], Faran [15], and Huang [20], every proper map from $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ for $N\leq 2n-2$ that is sufficiently smooth up to the boundary is spherically equivalent to the linear embedding.
+For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding.
+Interestingly, this map has no analogue for annuli.
+In general, see the books by D’Angelo [9, 11, 12] for a more thorough treatment of proper maps of balls.
+Let us move to maps of annuli.
+When $n=1$, the same classical argument to compute the automorphism group of $\mathbb{A}_{1,r}$ extends to show that a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case $f(z)=e^{i\theta}z^{d}$ or $f(z)=r^{d}e^{i\theta}z^{-d}$.
+As we said before, if $n>1$, annulus maps extend to ball maps.
+By Forstnerič, the maps must be rational.
+See Theorem˜2.2 (proved in [19]) for more details.
+So when $n>1$, holomorphic proper maps of annuli are in one-to-one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero.
+A natural map that is a proper map for both balls and annuli is the homogeneous map $z^{\otimes d}$ (the $d$th tensor power).
+That is because if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
+(1) $\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
 is constant.
-Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin.
-Moreover, it is not difficult to show that it takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N$ is the number of components of $H_{d}$ and $R=r^{d}$.
-Similarly, if we have any two proper maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$.
-We call such a weighted direct sum a juxtaposition.
-So juxtapositions of proper maps of annuli are proper maps of annuli as long as the source annulus is the same.
+This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary map to what we will call $H_{d}$ direct sum a number of zero components.
+In particular, $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}$.
+Rudin [23] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$.
+See D’Angelo [8, 10, 9] for a much simpler proof of this fact and many more details and results about the homogeneous map.
+Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin, and hence takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N=\binom{n+d-1}{d}$ is the number of components of $H_{d}$ and $R=r^{d}$.
+We can now construct new maps by what we will call juxtaposition.
+Given two proper annulus maps with the same source, $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$.
 Based on the one-dimensional result, noting that $z^{\otimes d}$ is the correct multidimensional analogue of $z^{d}$, one could (incorrectly) conjecture that all such proper maps are juxtapositions of homogeneous maps.
-However, other maps do exist, see Example˜2 and more generally [helal-2025-proper].
-Our goal is to start a classification of proper maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$.
+However, other maps do exist, see Section˜6 and more generally [19].
+Our goal is to classify proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$.
 We say $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ and $g\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are unitarily equivalent (or simply equivalent) if there exist unitary matrices $V\in U(n)$ and $W\in U(N)$ so that
@@
 As we noted, $U(n)$ is the holomorphic automorphism group of $\mathbb{A}_{n,r}$.
-If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.6.
+If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.7.
 If $N>n$, then an affine embedding exists whenever $r\leq R<1$, but is not necessarily $z\oplus 0$ due to the different automorphism group.
-In particular, $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
+In particular, if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
 (3) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0,$
-where the last 0 has Nn1N-n-1 terms. That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component. See Proposition˜2.7.
-As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney mapping exists and is not spherically equivalent to a linear embedding.
+where the last 0 has Nn1N-n-1 terms.
+That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component.
+Allowing for constants in juxtaposition, this map is a juxtaposition of the identity $z$ and the constant map $1$.
+See Proposition˜2.8.
+As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney map exists and is not spherically equivalent to a linear embedding.
 Curiously, for annulus maps, this gap is much longer and goes until $N=\binom{n+1}{2}$.
@@
 Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $2\leq n<N<\binom{n+1}{2}$.
-Then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$.
+Then $r\leq R<1$ and $f$ is unitarily equivalent to an affine embedding (3).
 The bound on $N$ is sharp as the homogeneous map $H_{2}$ takes the annulus $\mathbb{A}_{n,r}$ properly to $\mathbb{A}_{\binom{n+1}{2},r^{2}}$.
-As homogeneous maps play an outsize role in the study of proper maps of annuli we next find a property that determines a map to be homogeneous.
-The framework that we find applies more broadly to proper maps of balls and so we state it first in that setting.
-First we define a certain invariant for proper mappings that has, in various guises, appeared in the study of ball maps.
-First let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any holomorphic map.
+As homogeneous maps play a significant role in the study of proper maps of annuli, we next find a property that determines a map to be homogeneous.
+We define an invariant $k_{f}$ that has appeared in many guises in the study of ball maps.
+Let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any map.
 If $X\subset\mathbb{C}^{N}$, write $\operatorname{aff}(X)$ for the affine hull of $X$, that is, the smallest affine subspace of $\mathbb{C}^{N}$ containing $X$.
@@
 One could think of $k_{f}$ as a certain measure of complexity.
-It is not difficult to see that $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps.
-While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls has $k_{f}\leq N-1$.
-The idea that Dor [dor-1990-proper] used to show a proper map is not rational is to, in other words, show that the constructed map has $k_{f}=N$ as any rational ball map has $k_{f}=N-1$.
-If $k_{f}=N-1$, then we obtain a map of Grassmannians.
-That technique was used for example by Faran to prove his classification of maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$.
-The number $k_{f}$ is related to the $X$-variety that Forstnerič used to show the rationality result.
-In particular, for a rational map, if the dimension of the generic fiber of the $X$-variety is $\ell$, then $k_{f}=N-\ell-1$.
-D’Angelo [dangelo-2003-Xvariety] has studied the $X$-variety for rational maps and developed a technique of quickly computing fibers of the $X$-variety as the nullity of a certain matrix $C(w)$.
-The nullity of $C(w)$ for a generic $w$ is then the $\ell$.
-A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{B}_{n})$ (or $f(\mathbb{A}_{n,r})$).
-In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{B}_{n})$.
-In either case, restricting to $S$ on the target we obtain a proper map between balls or annuli with target dimension $N_{f}$.
-While for rational proper maps, often $k_{f}=N_{f}-1$, it is not always true.
-Example˜2 mentioned above has $k_{f}=N_{f}-3$.
-A juxtaposition of two rational maps also has $k_{f}\leq N_{f}-2$, see Proposition˜2.12.
+For a holomorphic map $f$, $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps.
+While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls (or annuli) has $k_{f}\leq N-1$.
+This fact has been used many times in the study of ball maps.
+The integer $k_{f}$ is also related to the $X$-variety.
+See Section˜3 for more on this connection and the history.
+A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{A}_{n,r})$ (or $f(\mathbb{B}_{n})$).
+In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{A}_{n,r})$.
+In either case, restricting the codomain to $S$ obtains a proper map between annuli (or balls) with target dimension $N_{f}$.
+Note that it is not always true that $k_{f}=N_{f}-1$.
+If $f$ is a juxtaposition of two rational maps, then $k_{f}\leq N_{f}-2$, see Proposition˜3.5.
 What we prove is that for proper maps of annuli, $k_{f}=N_{f}-1$ is precisely the condition that characterizes homogeneous maps $H_{d}$.
 Theorem 1.2.
-Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
+Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
 (5) $f=UH_{d},$
 for some $d\in\mathbb{N}$ and a unitary $U\in U(N)$.
-One direction of this theorem is obvious, it is that $k_{f}=N-1$ implies that $f$ is homogeneous that is remarkable.
-The proof relies on a result from [helal-2025-proper], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps.
-We note that if we wish to state the theorem in terms of $N_{f}$, then we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$.
-We note that $k_{f}=N-1$ implies that $N=N_{f}$ for any proper rational map of annuli (or balls).
+One direction of this theorem is obvious.
+The remarkable direction is that $k_{f}=N-1$ implies that $f$ is homogeneous.
+The proof relies on a result from [19], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps.
+To state the theorem in terms of $N_{f}$, we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$.
+Note that $k_{f}=N-1$ implies $N=N_{f}$ for any proper rational map of annuli (or balls).
 Let us state a version of the forward direction of this result for proper maps of balls.
@@
 Theorem 1.3.
-Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a rational proper map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$.
+Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a proper holomorphic map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$.
 Then there exists a unitary map $U$ and an integer $d$ such that $f=UH_{d}$.
-The proof follows by using the $k_{f}=N-1$ condition to show that $f$ must in fact take infinitely many spheres centered at zero to spheres centered at zero.
+To prove the theorem, we note that we have a rational map taking two distinct spheres to two distinct spheres.
+The condition $k_{f}=N-1$ is then used to show that $f$ must in fact take infinitely many spheres to spheres.
 As a corollary to Theorem˜1.2 and to the result of Faran classifying proper maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$, we completely classify proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$.
@@
     where $R=r^{2}$. Note that $N_{f}=3$ and $k_{f}=2$.
+It is also interesting to classify maps by degree.
+Degree one maps are unitaries possibly composed with an affine embedding if the target dimension is larger than the embedding dimension.
+We completely classify degree two maps in Section˜6.
+There exist other degree two maps than the homogeneous map $H_{2}$ or a juxtaposition of $H_{2}$ and $H_{1}$ and/or $H_{0}$.
+For proper ball maps, all degree two maps are spherically equivalent to monomial maps by the work of the second author [21].
+For annuli maps where the automorphism group is the unitary, the conclusion is more elaborate.
+For every $r$ and $R$ with $r^{2}<R<1$, there exists a one-parameter family of rational proper maps from $\mathbb{A}_{n,r}$ to $\mathbb{A}_{N,R}$, where $N_{f}=N=\binom{n+1}{2}+n$ and these maps are not equivalent to a polynomial map.
 Let us give the organization of this paper.
-In Section˜2, we state and prove the basic results for maps of annuli.
-In Section˜3, we prove the first gap theorem, that is, Theorem˜1.1.
-In Section˜4, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2.
+In Sections˜2 and 3, we state and prove the basic results for maps of annuli and the general hyperplane rank $k_{f}$.
+In Section˜4, we prove the first gap theorem, that is, Theorem˜1.1.
+In Section˜5, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2.
 We also prove the classification of annuli maps in dimension 2 and 3, that is, Theorem˜1.4.
-Finally, in Section˜5, we give some further remarks and open questions.
+In Section˜6, we give a complete classification of second degree rational annuli maps.
+Finally, in Section˜7, we give some further remarks and open questions.
+The authors would like to acknowledge John D’Angelo for many insightful comments on this work, and the second author in particular would like to acknowledge D’Angelo for introducing him to the problems surrounding proper maps of balls and for the many years of extremely fruitful discussions on related topics.
+The authors would also like to acknowledge Dusty Grundmeier for useful suggestions on this work.
 2. Basic results on proper maps of annuli
-Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper mappings of annuli in one dimension.
+Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper maps of annuli in one dimension.
 Proposition 2.1.
@@
 Let us sketch the proof as it may not be easy to find, but it is almost identical to classifying biholomorphisms of annuli.
-First using an inversion and scaling and the maximum principle, we may assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner.
-The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and goes to 0 on the outer circle and 1 on the inner circle.
+After a possible inversion and scaling, we assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner circle.
+The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and is 0 on the outer circle and 1 on the inner circle.
 Hence it is equal to $\frac{\log\mathopen{}\mathclose{{\left\lvert z}}\right\rvert}{\log r}$, and we get $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{\frac{\log R}{\log r}}$.
@@
 ∎
-In [helal-2025-proper], the authors have proved, using the theorem of Forstnerič [forstneric-1989-extending] as a key ingredient, that all annulus maps are rational when $n\geq 2$.
-For completeness we reproduce it here.
-Theorem 2.2 ([helal-2025-proper]).
+In [19], the authors have proved, using the theorem of Forstnerič [17] as a key ingredient, that all annulus maps are rational when $n\geq 2$. For completeness we reproduce it here.
+Theorem 2.2 ([19]).
 Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map.
@@
 Unlike in the ball case, no hypothesis on the boundary regularity is needed due to the pseudoconcavity of the inner boundary.
-The best expression for $D$ is not known, though D’Angelo has conjectured [dangelo-2003-sharp] that it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 is $n=2$.
+The best expression for $D(n,N)$ is not known, though D’Angelo has conjectured [3] that for ball maps, it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 if $n=2$.
+For annulus maps, we know that the $D$ must be smaller.
+For example, by our result for $n=2$ and $N=3$, we see that $D(2,3)=2$ in the annulus case, while Faran shows that in the ball case the degree bound is $3$.
+We conjecture that for annulus maps in dimension $n\geq 2$ and degree $d$, we have $N\geq\binom{n+d-1}{d}$.
 Sketch of proof.
@@
 ∎
-Via a theorem of Cima–Suffridge [cima-1990-boundary], a rational proper map of balls has no poles on the spheres, so we get a sphere map.
-Conversely a rational map taking a sphere to a sphere is a proper ball map via maximum principle.
-In [helal-2025-proper], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere.
-Putting all this together one gets the following equivalence.
+Via a theorem of Cima–Suffridge [2], a rational proper map of balls has no poles on the spheres, so we get a sphere map.
+Conversely, a rational map taking a sphere to a sphere is a proper ball map via maximum principle.
+In [19], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere.
+Putting all these together one gets the following equivalence.
 Corollary 2.3.
@@
 First, we construct a nonrational proper map from the disc $\mathbb{D}$ to the 2-ball $\mathbb{B}_{2}$.
-Such a map follows from the result of Dor [dor-1990-proper], but it is useful to consider an elementary construction in the one-dimensional case.
+Such a map follows from the result of Dor [13], but it is useful to consider an elementary construction in the one-dimensional case.
 Take an arbitrary nonrational function from the disc $f_{1}\colon\mathbb{D}\to\mathbb{D}$, suppose that it extends continuously to the closed disc $\overline{\mathbb{D}}$ but does not extend holomorphically past any point on $\partial\mathbb{D}$ and suppose that $\sup_{z\in\overline{\mathbb{D}}}\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert<1$.
 Find a real-valued harmonic function $u$ on $\mathbb{D}$ continuous on $\overline{\mathbb{D}}$ so that on $\partial\mathbb{D}$ we have $u(z)=\frac{1}{2}\log\mathopen{}\mathclose{{\left(1-\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}}}\right)$.
-Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$.
-Then write $f_{2}=e^{u+iv}$.
-Then the mapping $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$.
-Clearly $f$ is not spherically equivalent to a rational map if $f_{1}$ is not rational.
+Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$ and write $f_{2}=e^{u+iv}$.
+Then the map $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$.
+If $f_{1}$ is not rational, then $f$ is not spherically equivalent to a rational map.
 The map $f$ is at most points locally an embedding.
@@
 Let us compare the effect of the automorphism groups for proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ versus $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$.
-The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form (9) below.
+The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form
+(8) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z.$
 For any affine subspace $L$ of dimension $m<n$ such that $L\cap\mathbb{B}_{n}\not$$\varnothing$, there exists a $\varphi\in\operatorname{Aut}(\mathbb{B}_{n})$ so that $\varphi(L)$ is given by $z_{m+1}=\cdots=z_{n}=0$.
@@
 For a proper ball map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ where $N_{f}<N$, there exists an automorphism $\psi\in\operatorname{Aut}(\mathbb{B}_{N})$ so that $\psi\circ f=F\oplus 0$ where $F\colon\mathbb{B}_{n}\to\mathbb{B}_{N_{f}}$ is a proper map ($N_{f}=N_{F}$).
-The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space, it can only “rotate” it.
+The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space; it can only “rotate” it.
 The proof is an elementary computation.
@@
 Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N_{f}<N$. Then there exists a unitary $U\in U(N)$ and a unique $c\in[0,R)$ such that
-(8) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0$
+(9) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0,$
 where $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{f},\sqrt{\frac{R^{2}-c^{2}}{1-c^{2}}}}$ is a proper map and $N_{f}=N_{F}$.
 In the proposition, the 0 in the direct sum has $N-N_{f}-1$ components (so it can be left out if $N=N_{f}+1$).
-The proposition leads to classification of first degree proper rational maps.
-We first start with the equidimensional case.
-In this case, unitaries (so unitarily equivalent to the identity) are the only maps.
+To classify annulus maps the following division lemma is very useful.
+If $f=\frac{p}{q}$ is a rational proper annulus map, then it takes the unit sphere to the unit sphere and the $r$-sphere to the $R$-sphere.
+Namely, $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}$ is divisible by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1$ and $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-R^{2}\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}$ is divisible by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}$.
+The following lemma encapsulates both of these conditions.
+The lemma is a special case of Lemma 3.11 from [19].
+Lemma 2.5.
+Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree $d$, $n\geq 2$, and write $f=\frac{p}{q}$ in lowest terms. Then $q$ is of degree at most d1d-1 and
+(10) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\Bigr)\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
+for a real polynomial $Q(z,\bar{z})$ and $b=\frac{1-R^{2}}{1-r^{2}}$. The $Q$ is of degree at most d2d-2 in $z$ and d2d-2 in $\bar{z}$.
+Let us use the lemma to classify maps of degree 1.
+We consider the equidimensional case.
+In the proof we will use the notion of a rank of a real polynomial $\rho(z,\bar{z})$ in $\mathbb{C}^{n}$.
+By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
+(11) $\rho(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
+where $\epsilon_{j}=\pm 1$ and $P_{j}$ are holomorphic polynomials.
 Proposition 2.6.
@@
 Proof.
-We have that $f$ is an LFT that must take the unit sphere to the unit sphere, hence it is well-known that it must have the form
-(9) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z,$
-where $U$ is a unitary and $a\in\mathbb{B}_{n}$. We need to show that $a=0$. By precomposing with a unitary, we can assume that $a=(a_{1},0,\ldots,0)$ and
-(10) $\phi_{a}(z)=\mathopen{}\mathclose{{\left(\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}},\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{2}}{1-z_{1}\bar{a}_{1}},\ldots,\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{n}}{1-z_{1}\bar{a}_{1}}}}\right).$
-We note that the complex line $z_{2}=\cdots=z_{n}=0$ gets taken by $\phi_{a}$ to this same complex line, and on this line $\phi_{a}$ is simply the corresponding LFT in one dimension $\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$.
-It remains to show that if $z_{1}\mapsto\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$ takes a circle of radius $r<1$ centered at zero to a circle of radius $R<1$ centered at zero.
-In other words, we must have that for all $\theta$, $\mathopen{}\mathclose{{\left\lvert a_{1}-re^{i\theta}}}\right\rvert^{2}=R^{2}\mathopen{}\mathclose{{\left\lvert 1-re^{i\theta}\bar{a}_{1}}}\right\rvert^{2}$.
-By expanding and equating coefficients one obtains that $a_{1}=0$ and $r=R$.
-In other words, $a=0$ and $f$ is a unitary.
+We apply Lemma˜2.5 to find $b$ and $Q$.
+As $f$ is of degree 1, then $Q\equiv 0$ and $q$ is a constant, meaning we can just take $q=1$.
+The rank of $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ must be the target dimension $n$, but the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=(1-b)+b\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is $n+1$ unless $b=1$.
+So $b=1$ and $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$, which implies that $f(z)=Uz$ for a unitary $U$.
 ∎
-As a quick corollary we can remove the hypothesis of rationality and degree.
-By Hartogs phenomenon and the maximum principle, we find that a holomorphic proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself.
+As a quick corollary, we can remove the hypothesis of rationality and degree.
+By Hartogs phenomenon and the maximum principle, we find that a proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself.
 That this map is a rational map of degree 1 then follows by Alexander’s theorem.
@@
 If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ is a proper holomorphic map and $n\geq 2$, then $r=R$ and $f$ is a unitary map.
-When $N>n$, the result is a little more complicated.
+When $N>n$, the result is a little more complicated but follows in the same way.
 Proposition 2.8.
@@
 (12) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0.$
-In the proof we will use the notion of a rank of a real polynomial $r(z,\bar{z})$ in $\mathbb{C}^{n}$.
-By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
-(12) $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
-where $\epsilon_{j}\pm 1$ and $P_{j}$ are holomorphic polynomials.
 Proof.
-Write $f=\frac{p}{q}$. As $f$ takes the unit sphere to the unit sphere we have
-(13) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert q(z)}}\right\rVert^{2}=\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$
-where $\lambda$ has to be a constant as $p$ and $q$ are of degree 1.
-Therefore, the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ has to be $n$ as that is the number of positive eigenvalues of the matrix of coefficients $\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$.
-In other words, $N_{f}=n$ and we can apply Proposition˜2.4.
-The $F$ in the proposition is unitarily equivalent to the identity by Proposition˜2.5.
-The values of the factors follow by elementary computation.
+Again apply Lemma˜2.5 to find $q=1$ and $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=(1-b)+b\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$.
+For this to be a squared norm we must have $b\leq 1$, which implies $R\geq r$.
+The form then follows as above.
 ∎
-We derive a few of basic results about the $k_{f}$ integer.
+3. General hyperplane rank and its connection to the $X$-variety
+Before proving some basic results about $k_{f}$ that we will require, let us discuss some of the history of this invariant.
+When Dor [13] constructed a nonrational map of balls, to show nonrationality, he showed that (in our language) $k_{f}=N$, as any rational ball map has $k_{f}\leq N-1$.
+When $k_{f}=N-1$, we obtain a map of Grassmannians that takes an open dense set of affine hyperplanes in $\mathbb{C}^{n}$ to affine hyperplanes in $\mathbb{C}^{N}$.
+This technique was used by Faran to classify proper maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$.
+The number $k_{f}$ is also related to the $X$-variety that Forstnerič used to show the rationality result.
+Let us quickly define the $X$-variety and note the connection for rational maps.
+Let $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ be a rational proper map.
+By complexifying the relation that $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=1$ whenever $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}=1$, we find $\langle f(z),f(w)\rangle=1$ whenever $\langle z,w\rangle=1$.
+The $X$-variety is defined as
+(13) $X_{f}=\{(z,\zeta)\in\mathbb{C}^{n}\times\mathbb{C}^{N}:\langle\zeta,f(w)\rangle=1\text{ for all $w$ such that }\langle z,w\rangle=1\}.$
+Note that the equation for (almost) every affine hyperplane in $\mathbb{C}^{n}$ can be written as $\langle z,w\rangle=1$ for a fixed $w$.
+Moreover, $\langle\zeta,f(w)\rangle=1$ gives affine hyperplanes on the target.
+The fiber of the $X$-variety is the subset for a fixed $z$.
+Such a fiber is an affine subspace of some dimension.
+It is not difficult to show that if the generic fiber of the $X$-variety has dimension $\ell$, then $k_{f}=N-\ell-1$.
+The $X$-variety encodes everything about the map $f$.
+For example, if $\ell=0$, then we find that the $X$-variety is (except some possible exceptional fibers) the graph of the map $f$.
+It also encodes the multivariable analogue of reflection across the sphere.
+The $X$-variety was defined by Forstnerič for the rationality result, and it has been extensively studied by D’Angelo, see [6, 7].
+In fact, due to the connection of $k_{f}$ and the dimension of the fiber, some of the basic results about the $k_{f}$ integer follow directly from the work of D’Angelo, although we present direct proofs here.
+In particular D’Angelo developed a technique of quickly computing the fibers of the $X$-variety by associating to the map a certain matrix $C(w)$ whose nullspace gives the fiber of the $X$-variety at each point.
+Bounding the general hyperplane rank using Green’s hyperplane restriction theorem from commutative algebra was also used in a paper by Grundmeier, the second author, and Vivas [18] to prove a rigidity theorem for CR maps between hyperquadrics.
+Let us then derive a few basic results about the $k_{f}$ integer.
 First we note that for any rational map, it is enough to check an open dense subset of hyperplanes, that is, the general hyperplane.
-Similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication.
+A similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication.
 The space of affine hyperplanes in $\mathbb{C}^{n}$ has the topology of the Grassmannian (minus the hyperplane at infinity).
@@
 Proposition 3.2.
-If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
+Suppose $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
 Proof.
@@
 Writing $P=q^{-1}(0)$ for the pole set of $f$, we therefore have that $f(\Sigma_{w}\setminus P)\subset\Sigma_{f(w)}$ for all $w$ where $f(w)$ is defined.
-As any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$.
+Any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$.
 Therefore, there is an open set of affine hyperplanes which $f$ takes into hyperplanes.
@@
 ∎
-The $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
+The general hyperplane rank $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
 Proposition 3.3.
@@
 Next we consider the $k_{f}$ integer for juxtapositions.
-We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [dangelo-2009-complexity].
+We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [4].
 The next proposition therefore shows that for rational ball (or annulus) maps, $k_{f}=N_{f}-1$ happens only on the non-extremal points of this convex set.
@@
 Suppose that $F_{t}$ is a juxtaposition of two rational proper maps of balls (or annuli) $F_{0}$ and $F_{1}$, that is, $F_{t}=\sqrt{1-t}\,F_{0}\oplus\sqrt{t}\,F_{1}$ for some $t\in(0,1)$, and further suppose that $\mathopen{}\mathclose{{\left\lVert F_{0}(z)}}\right\rVert^{2}\not\equiv\mathopen{}\mathclose{{\left\lVert F_{1}(z)}}\right\rVert^{2}$.
-Then $k_{f}\leq N_{f}-2$.
+Then $k_{F_{t}}\leq N_{F_{t}}-2$.
 The condition on the norms is to ensure that we are taking the convex combination of a point with itself in the space of squared norms.
@@
 Therefore suppose that $k_{F_{0}}=N_{F_{0}}-1$ and $k_{F_{1}}=N_{F_{1}}-1$.
-Also note from the proof of Proposition˜2.9 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies.
-In particular, we find these vectors in an $N_{f}$-dimensional hyperplane which we can assume is the entire target space.
-Apart from these facts that we can simply treat these as rational maps.
+Also note from the proof of Proposition˜3.2 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies.
+In particular, we find these vectors in an $N_{f}$-dimensional hyperplane, which we can assume is the entire target space.
 The integers $k_{f}$ and $N_{f}$ and the fact about the normal vectors are all invariant under any LFT and so we only need to prove the result for $F_{0}\oplus F_{1}$.
@@
 The dimensions increase by 1, but that is not relevant to us as we are interested in their difference.
-In other words, what we have is two homogeneous polynomial maps of the same degree $F_{0}$ and $F_{1}$, we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent.
+In other words, what we have is two homogeneous polynomial maps $F_{0}$ and $F_{1}$ of the same degree, and we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent.
 We further have that for any codimension 1 subspace $H$, the span of $F_{0}(H)$ is of codimension 1 (in the codomain of $F_{0}$) and same for $F_{1}(H)$.
 It is possible that components of $F_{0}$ and $F_{1}$ are not linearly independent when put together.
-After composing with a linear we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components.
+After composing with a linear map, we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components.
 Write the target space as $\mathbb{C}^{j}\times\mathbb{C}^{k}\times\mathbb{C}^{\ell}$.
-The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ ($v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, $0\in\mathbb{C}^{\ell}$) so that $\xi\cdot v=0$ for all $\xi\in P(H)$.
+The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ with $v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, and $0\in\mathbb{C}^{\ell}$ so that $\xi\cdot v=0$ for all $\xi\in P(H)$.
 There is also a vector $w=(w_{1},0,w_{3})$ so that $\xi\cdot w=0$ for all $\xi\in P(H)$.
-We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, we found that $P(H)$ lives in a codimension 2 space and we are done.
+We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, $P(H)$ lives in a codimension 2 space and we are done.
 Thus suppose that there exists a $v=w=(v_{1},0,0)$ for every subspace $H$.
@@
 ∎
-We have seen that the homogeneous map $H_{d}$ is an annulus map.
-We also have that juxtapositions of homogeneous maps and their affine embeddings are annulus maps.
-Here is an example that is not constructed in this way from homogeneous maps.
-Example 2.
-With the construction introduced in [helal-2025-proper], one finds the following map
-(16) $f(z_{1},z_{2})=\mathopen{}\mathclose{{\left(\tfrac{2}{\sqrt{5}}z_{1}^{3},\tfrac{2\sqrt{2}}{\sqrt{5}}z_{1}^{2}z_{2},\tfrac{2}{\sqrt{5}}z_{1}z_{2}^{2},z_{1}z_{2},z_{2}^{2},\tfrac{1}{\sqrt{5}}z_{1}}}\right).$
-This map gives a proper map of $\mathbb{A}_{2,\frac{1}{2}}$ to $\mathbb{A}_{6,\frac{1}{4}}$.
-It takes the unit sphere to the unit sphere and the half sphere to quarter sphere.
-It takes no other sphere centered at the origin to a sphere centered at the origin.
-It is not unitarily equivalent to a juxtaposition of several homogeneous maps, nor is it an affine embedding of such a map.
-If it were, then it would take all spheres centered at the origin to spheres centered at the origin.
-We have that $N_{f}=6$ and one can compute that the $k_{f}$ number is $3$, so $k_{f}<N_{f}-1$.
-To see why it is $3$, note that if we restrict to a 1-dimensional affine space $H$, the map becomes a one variable cubic polynomial and so the affine span can be at most $3$.
-That it is exactly $3$ can be seen by restricting to a hyperplane making $z_{2}$ constant.
-We remark that while Theorem˜1.2 implies that for annuli maps, $k_{f}=N-1$ means $f$ is homogeneous, there must be extremal points with $k_{f}<N-1$ in the convex set describing the rational proper annulus maps as the example above shows that there exist maps that are not constructed out of homogeneous maps via juxtapositions.
-The construction from [helal-2025-proper] can be used to obtain cubic examples in higher dimensions and also quadratic rational maps with the same properties (not taking more than two zero-centric spheres to zero-centeric spheres, so not constructed out of homogeneous maps).
 4. First gap
-Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [dangelo-2012-pfisters]*Theorem 1.1.
+Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [5]*Theorem 1.1.
 We only need the quadratic version of this result, but we state the more general result.
@@
 (16) $\operatorname{rank}Q(z,\bar{z})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})\geq\operatorname{rank}(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})=\binom{n+2}{2}$.
-The proof follows by roughly the same argument as in [dangelo-2012-pfisters] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors.
+The proof follows by roughly the same argument as in [5] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors.
 What must be newly proved here is a “diagonal version” of the theorem.
@@
 Proof.
-The main new claim that must be proved is the “diagonal version” of this theorem. That is, for a real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
+The main new claim that must be proved is the “diagonal version” of this theorem. That is, for real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
 (17) $\bigl(Q(x)(x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)\\ \geq\#\bigl((x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)=\binom{n+2}{2}$,
 where #\# means the number of nonzero coefficients in the given polynomials when they are expanded.
-We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{d}$ and applying the claim for positive $s_{1},s_{2}$.
+We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{2}$ and applying the claim for positive $s_{1},s_{2}$.
 First suppose that $n=1$. For this case, assume $s_{1}$ and $s_{2}$ are both negative. The polynomial
@@
 Write $R(x)=x_{1}^{j}x_{2}^{k}\widetilde{R}(x)$ (and similarly $Q(x)=x_{1}^{j}x_{2}^{k}\widetilde{Q}(x)$) where $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ are both not identically zero.
-Applying, the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively).
+Applying the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively).
 One of these terms can be the constant, and therefore, we have $1+2n$ terms in $\widetilde{R}(x)$ which are all pure powers.
@@
 As $\widetilde{Q}$ is a polynomial, $L^{\nu}\widetilde{Q}$ must be identically zero for some $\nu$, so take $\nu$ such that $L^{\nu-1}Q$ is not zero but $L^{\nu}Q\equiv 0$.
-Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation and so it is a function $x_{1}+x_{2}$, that is $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$ $S$ is not identically zero.
+Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation $Lu=0$, and so it is a function of $x_{1}+x_{2}$, that is, $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$, $S$ is not identically zero.
 But then
@@
 Since $L^{\nu-1}\widetilde{R}(x)$ has a mixed monomial, then so does $\widetilde{R}(x)$ as $L$ cannot manufacture mixed monomials out of pure ones.
-Therefore, $R(x)$ has at least $1+2\cdot 2+1=\binom{2+2}{2}=6$ terms.
-Moreover, we guarantee that there are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$, and at least 3 terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m\geq k$.
-Note that also there are no terms at all with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$.
-Similarly, there are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$, and at least 3 terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell>k$ and $m\geq j$.
-And no terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell<k$.
+Therefore, $R(x)$ has at least $1+2\cdot 2+1=6=\binom{2+2}{2}$ terms.
+In particular, we have the following terms:
+Statement 1.
+There are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$.
+There are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$.
+There is at least one term with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m>k$.
+There are no terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$ or $m<k$.
 We now attack the general problem for $n>2$.
@@
 All the monomials of $r_{\mu}(x_{1},x_{2})(x^{\prime})^{\mu}$ must appear in $R(x)$ by minimality of $\mu$.
-We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these $a_{\ell m}$ is as in the “Moreover” remark above.
-We use this this remark for every pair of variables to count monomials.
+We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these monomials is given in ˜1.
+We use this result for every pair of variables to count monomials.
 For every pair there is one monomial that we might be also counting for other pairs, so we count it only once.
@@
 distinct monomials in $R(x)$. The claim is proved.
-Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [dangelo-2012-pfisters].
+Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [5].
 That is, there exists some ordering of monomials so that if we write out the matrix of coefficients of the expressions involved, multiplication by $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ moves along diagonals.
@@
 Up to multiplication with a monomial, the terms on this diagonal are polynomials in $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$, and so by replacing $x_{j}=\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ we obtain the situation of the claim and the proof follows.
-See [dangelo-2012-pfisters] for more details.
+See [5] for more details.
 ∎
@@
 Proof of Theorem˜1.1.
-By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$. Lemma 3.11 from [helal-2025-proper] gives us
-(25) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\Bigr)\cdot\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q_{2}(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
-for a real polynomial $Q_{2}(z,\bar{z})$.
-If $Q_{2}\equiv 0$, then by the argument in the proof of Theorem 3.12 from [helal-2025-proper], $f$ is a polynomial, that is, $q\equiv 1$.
-Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$).
+By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$ in lowest terms. By Lemma˜2.5 we have
+(24) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\Bigr)\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
+for a real polynomial $Q(z,\bar{z})$ and $b=\frac{1-R^{2}}{1-r^{2}}$.
+If $Q\equiv 0$, then since $\frac{p}{q}$ is in lowest terms, we must have $q$ is constant, so we may assume $q\equiv 1$.
+Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$).
 Therefore, $\deg f=1$.
-Proposition˜2.7 implies that $f$ is an affine embedding and that $R\geq r$.
-Suppose $Q_{2}\not\equiv 0$, we will get a contradiction.
+Proposition˜2.8 implies that $f$ is an affine embedding and that $R\geq r$.
+Suppose $Q\not\equiv 0$, we will get a contradiction.
 By applying Proposition˜2.4, we reduce to the case when $N=N_{f}$.
@@
 Therefore, $N$ is equal to the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$.
-Theorem˜3.1 gives us
+Theorem˜4.1 gives us
 […]
@@
 We therefore obtained $N_{f}=n$ and that the degree is $1$.
-The conclusion follows from Proposition˜2.5 or Proposition˜2.7 depending on whether $N>n$ or $N=n$.
-The interval is sharp; when $N=\binom{n+1}{2}$, then besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding.
+The conclusion follows from Proposition˜2.6 or Proposition˜2.8 depending on whether $N>n$ or $N=n$.
+The interval is sharp; when $N=\binom{n+1}{2}$, besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding.
 ∎
@@
 See Theorem˜1.4.
-4. Homogeneous maps via the generic hyperplane rank
+5. Homogeneous maps via the general hyperplane rank
 To prove the theorem on classification of homogeneous maps, we will prove a more general result about rational maps taking two spheres to spheres.
@@
 Our first result is about rational maps taking two LFT spheres to LFT spheres.
-The definition of $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$.
+The definition of the general hyperplane rank $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$.
 In the proof we will work in homogeneous coordinates on the projective space, and so a general hyperplane in $H\subset\mathbb{P}^{n}$ is an $n$-dimensional linear subspace in $\mathbb{C}^{n+1}$.
 We use $z=(z_{0},\ldots,z_{n})$ for the homogeneous coordinates on $\mathbb{P}^{n}$.
-To write a rational function $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
+To write a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
 The sphere $S^{2n-1}\subset\mathbb{C}^{n}\subset\mathbb{P}^{n}$ in our chosen affine embedding when put into homogeneous coordinates is given by the equation
@@
 Theorem 5.1.
-Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to two LFT spheres and $k_{f}=N-1$.
+Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to LFT spheres and $k_{f}=N-1$.
 Then $f\circ\alpha=\beta\circ f$ for some LFTs $\alpha\in\operatorname{Aut}(\mathbb{P}^{n})$ and $\beta\in\operatorname{Aut}(\mathbb{P}^{N})$ neither of which is the identity.
 More precisely, using the coordinates above, if one of the spheres on the source (resp.
-the target) are given by $J_{n}$ (resp.
+the target) is given by $J_{n}$ (resp.
 $J_{N}$) and the other by a hermitian matrix $A$ (resp.
@@
 To figure out the map one uses the Segre variety of the LFT spheres.
-But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different mapping of Grassmannians.
+But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different map of Grassmannians.
 Comparing the two maps gives us the corresponding $\alpha$ and $\beta$.
@@
 As we said we will work in homogeneous coordinates, that is, in $\mathbb{C}^{n+1}$ and $\mathbb{C}^{N+1}$.
+We will write $z$ and $w$ for points in $\mathbb{C}^{n+1}$ and $Z$ and $W$ for points in $\mathbb{C}^{N+1}$.
 Starting with a rational map $f$, we clear denominators and homogenize to represent $f$ by a homogeneous polynomial map $F=(F_{0},F_{1},\ldots,F_{N})\colon\mathbb{C}^{n+1}\to\mathbb{C}^{N+1}$ where $F_{0}$ is the denominator.
-After applying an LFT we can assume that one of the matrices on both the source and the target is the standard sphere $S^{2n-1}$ and $S^{2N-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$.
+After applying an LFT we can assume that one of the matrices on the source corresponds to the standard sphere $S^{2n-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$, and similarly for the target.
 For any hermitian matrix $M$ and a point $w$, we write the Segre variety of the homogenized LFT sphere given by $\mathopen{}\mathclose{{\left\langle Mz,z}}\right\rangle=0$ as
 (28) $\Sigma_{w}^{M}=\{z:\mathopen{}\mathclose{{\left\langle Mz,w}}\right\rangle=0\}.$
-In affine coordinates, the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z),p(w)}}\right\rangle-q(z)\overline{q(w)}=g(z,\bar{w})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right)$ for a polynomial $g$.
+Write $z=(z_{0},z^{\prime})$ where $z^{\prime}$ refers to the affine coordinates.
+Then the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z^{\prime}),p(w^{\prime})}}\right\rangle-q(z^{\prime})\overline{q(w^{\prime})}=g(z^{\prime},\bar{w}^{\prime})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z^{\prime},w^{\prime}}}\right\rangle-1}}\right)$ for a polynomial $g$.
 In homogeneous coordinates, this translates to
@@
 (35) $F(Sw)=\lambda J_{N}^{-*}B^{*}F(w).$
-Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$, we have
+Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$,
 (36) $F(Sw)=\lambda TF(w).$
@@
 It turns out that these spheres need not be the ones we started with.
-By iterating, we find that for any $k\in\mathbb{Z}$ we have that
+By iterating, we find that for any $k\in\mathbb{Z}$,
 (38) $f\circ\alpha^{k}=\beta^{k}\circ f.$
@@
 In our particular case, when all spheres involved are centered at zero, $f$ takes infinitely many spheres centered at zero of various radii to spheres centered at zero.
-In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, then we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
+In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
 Lemma 5.2.
@@
 Suppose also that $k_{f}=N-1$.
-Then there are two sequences of distinct numbers $\{r_{k}\}$ and $\{R_{k}\}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{K}S^{2N-1}$ for every $k$.
+Then there are two sequences of distinct numbers $\{r_{k}\}_{k=1}^{\infty}$ and $\{R_{k}\}_{k=1}^{\infty}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{k}S^{2N-1}$ for every $k$.
 Proof.
-We set things up as in the proof of Theorem˜4.1, and so the matrices $A$ and $B$ are given as
+We set things up as in the proof of Theorem˜5.1, and so the matrices $A$ and $B$ are given as
 (39) $A=\begin{bmatrix}-r^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}\quad\text{and}\quad B=\begin{bmatrix}-R^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
@@
 ∎
-We are now ready to prove Theorem˜1.2 using the results of our previous work in [helal-2025-proper].
+We are now ready to prove Theorem˜1.2 using the results of our previous work in [19].
 Proof of Theorem˜1.2.
 One direction of the theorem is clear and follows from the discussion so far.
-What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$.
-By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜4.2.
+What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$.
+By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜5.2.
 Therefore, $f$ takes infinitely many spheres centered at zero to spheres centered at zero.
-In [helal-2025-proper], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros.
+In [19], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros.
 So for some unitary $U$, we can assume that
-(42) $f=U(c_{1}H_{d_{1}}\oplus\cdots c_{\ell}H_{d_{\ell}}\oplus 0),$
+(41) $f=U(c_{1}H_{d_{1}}\oplus\cdots\oplus c_{\ell}H_{d_{\ell}}\oplus 0),$
 where we can assume that $d_{1},\ldots,d_{\ell}$ are distinct.
 Since we are assuming that $k_{f}=N-1$, we must have that $N_{f}=N$ and hence there are no zero components.
-Then, via the argument in the proof of Proposition˜2.11, we can see that $k_{f}=N$-$\ell$.
+Then, via the argument in the proof of Proposition˜3.4, we can see that $k_{f}=N$-$\ell$.
 Hence $\ell=1$, and $f=UH_{d}$ as required.
@@
 We already have that proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$ are rational.
-If the degree of the map is 1, then via Proposition˜2.7, it is unitarily equivalent to the affine embedding
+If the degree of the map is 1, then via Proposition˜2.8, it is unitarily equivalent to the affine embedding
 (42) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{1},\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{2},\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}}}\right),$
@@
 The theorem will therefore follow if we can show that any map of degree greater than 1 is unitarily equivalent to $H_{2}$.
-This result follows by Faran’s [faran-1982-maps] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$.
+This result follows by Faran’s [14] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$.
 The following lemma finishes the proof of Theorem˜1.4.
@@
 ∎
+6. Complete normal form of degree 2 annulus maps
+In the following theorem, it will be useful to write $H^{\prime}_{2}(z_{2},\ldots,z_{n})$, that is, the symmetrized homogeneous map in the variables $z_{2},\ldots,z_{n}$:
+(43) $H^{\prime}_{2}(z_{2},\ldots,z_{n})=\bigl(z_{2}^{2},\sqrt{2}z_{2}z_{3},\ldots,\sqrt{2}z_{2}z_{n},z_{3}^{2},\sqrt{2}z_{3}z_{4},\ldots,\sqrt{2}z_{3}z_{n},\ldots,z_{n}^{2}\bigr).$
+Note that when $n=2$, which is the main case to keep in mind, $H^{\prime}_{2}(z_{2})=z_{2}^{2}$.
+The complete classification of rational degree 2 maps of annuli is then the following theorem.
+One should contrast this with the much simpler classification for quadratic proper maps of balls, where up to spherical equivalence, every such map is monomial.
+All annuli maps are also ball maps, and hence all of the following maps are in fact spherically equivalent to monomial maps, but here we have to classify up to the unitary group.
+On the other hand, compared to the ball case, thanks to Lemma˜2.5, the classification of annuli degree 2 maps follows by quite tedious but elementary methods, not requiring sophisticated results in linear algebra.
+In order to simplify the statement we assume that $N=N_{f}$.
+Therefore, to get all degree two maps in general one would also have to allow compositions with linear embeddings.
+Theorem 6.1.
+Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 2, $n\geq 2$, such that $N_{f}=N$, and write $b=\frac{1-R^{2}}{1-r^{2}}$. Then $R\geq r^{2}$ and:
+  1. (i)
+    If $R=r^{2}$, then $N=\binom{n+1}{2}$ and $f$ is unitarily equivalent to $H_{2}$.
+  2. (ii)
+    If $r^{2}<R<r$ and $f$ is polynomial, then $N=\binom{n+1}{2}+n$ and $f$ is unitarily equivalent to the juxtaposition $\sqrt{1-t}\,H_{1}\oplus\sqrt{t}\,H_{2}$ for exactly one $t\in(0,1)$.
+  3. (iii)
+    If $r^{2}<R<r$ and $f$ is not polynomial, then $N=\binom{n+1}{2}+n$ and $f$ is unitarily equivalent to a map of the form
+    (44) $\begin{split}\Biggl(&\frac{\sqrt{1-b+Qr^{2}}+\frac{(1-b)a}{\sqrt{1-b+Qr^{2}}}z_{1}}{1+az_{1}},\frac{\sqrt{ba^{2}+Q}\,z_{1}^{2}+\frac{ba}{\sqrt{ba^{2}+Q}}z_{1}}{1+az_{1}},\\ &\qquad\frac{\sqrt{b-Q(1+r^{2})}+\frac{ba}{\sqrt{b-Q(1+r^{2})}}z_{1}}{1+az_{1}}\otimes(z_{2},\ldots,z_{n}),\\ &\qquad\frac{\sqrt{ba^{2}+2Q-\frac{b^{2}a^{2}}{ba^{2}+2Q}}\,z_{1}}{1+az_{1}}\otimes(z_{2},\ldots,z_{n}),\frac{\sqrt{Q}}{1+az_{1}}H^{\prime}_{2}(z_{2},\ldots,z_{n})\Biggr)\end{split}$
+    for exactly one pair $a$ and $Q$, where $a>0$,
+    (45) $a\leq\sqrt{\frac{br^{4}+(b^{2}-b+1)r^{2}-b+1-2r\sqrt{(b-1)b(1+r^{2})(br^{2}-b+1)}}{b^{2}r^{6}+2br^{4}+r^{2}}}$
+    and $Q$ is either one of
+    (46) $Q=\frac{-a^{2}br^{4}+(2(1-a^{2})b+a^{2}-1)r^{2}+b-1}{2(r^{4}+r^{2})}\\ \pm\frac{\sqrt{a^{4}r^{4}(br^{2}+1)^{2}-2a^{2}r^{2}(br^{2}(r^{2}+b-1)+1+r^{2}-b)+(1+r^{2}-b)^{2}}}{2(r^{4}+r^{2})}.$
+    Moreover, for each such pair $a$ and $Q$, a map exists.
+  4. (iv)
+    If $R=r$, then $N=\binom{n+1}{2}+n$ and $f$ is unitarily equivalent to a map of the form (44) for exactly one $a\in\mathopen{}\mathclose{{\left(0,\frac{1}{\sqrt{1+r^{2}}}}}\right)$ and $Q=\frac{1}{1+r^{2}}-a^{2}$.
+Moreover, for each such $a$, a map exists.
+  5. (v)
+    If $r<R<1$, then $N=\binom{n+1}{2}+n$ and $f$ is unitarily equivalent to a map of the form (44) for exactly one $a\in(0,1)$ where $Q$ is the larger value of (46).
+Moreover, for each such $a$, a map exists.
+We observe that by part (v) we obtain maps where the embedding dimension equals the target dimension and $R$ is arbitrarily close to $1$.
+It is not possible to do the same with juxtapositions of homogeneous maps only.
+The key part of the proof is to use Lemma˜2.5. In the case of degree 2 maps we quickly obtain the following first step in the classification.
+Proposition 6.2.
+Suppose $f=\frac{p}{q}\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 2, $n\geq 2$, take $q(0)=1$, and write $b=\frac{1-R^{2}}{1-r^{2}}$.
+Then after a possible unitary transformation on the source,
+(47) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\Bigr)\mathopen{}\mathclose{{\left\lvert 1+az_{1}}}\right\rvert^{2}+Q\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr),$
+where $0\leq a<1$ and $Q>0$ are constants and these two constants are the complete invariants for such maps under unitary equivalence.
+Proof.
+We apply Lemma˜2.5, and as the degree is 2, we find that $Q$ is constant and $q$ is of degree 1 or 0.
+We can take $q(0)=1$ and so, $q(z)=1+v\cdot z$ for some vector $v\in\mathbb{C}^{n}$.
+If $v=0$, we are in the polynomial case and $q\equiv 1$.
+Note that pre- or postcomposing with a unitary cannot change whether the map $f$ is polynomial or rational.
+In the rational case, that is, $v\not=0$, we precompose with a unitary so that $v\cdot z=az_{1}$ for some $a>0$.
+In the polynomial case, $v=0$ and hence $v\cdot z=az_{1}$ for $a=0$.
+In either case, precomposing $f$ (and hence $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$) with a unitary cannot change $a$ or $Q$ after this normalization.
+Postcomposing $f$ with a unitary is the same as postcomposing $p$ with a unitary and this leaves $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ unchanged.
+In other words, the normalized polynomial $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ is a complete invariant under unitary equivalence.
+It is not difficult to see that $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ uniquely determines $Q$ and $a$ and hence they are complete invariants under unitary equivalence.
+As $f$ is a rational ball map, we know that it extends past the boundary and hence $q$ cannot be zero on the closed unit ball, which means that $a<1$.
+By considering the second degree terms in $p$, that is, $\mathopen{}\mathclose{{\left\lVert p_{2}(z)}}\right\rVert^{2}=b\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}\mathopen{}\mathclose{{\left\lvert az_{1}}}\right\rvert^{2}+Q\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{4}$, we find that $Q\geq 0$ by setting $z_{1}=0$.
+Moreover, if $Q=0$, then every component in $p$ would be divisible by $q$, meaning that $f$ is not in fact a degree 2 map but a degree 1 map.
+So $Q>0$.
+∎
+Therefore to complete the proof of the theorem, we simply need to find conditions on $b$, $r$, $a$, and $Q$ so that the expression on the right hand side of ˜47 is a hermitian sum of squares.
+Writing $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ as a sum of hermitian squares is the same as writing the hermitian matrix of coefficients as a sum of outer products of vectors.
+These vectors are the coefficients of the polynomial components of $p(z)$.
+In particular, a real polynomial is a sum of hermitian squares if and only if its hermitian matrix of coefficients is a positive semidefinite matrix.
+If we are starting with the right hand side of ˜47 we can ensure that these vectors are linearly independent, which is equivalent to $f$ not mapping into a subspace.
+If the components of the map are linearly independent, the rank of this matrix of coefficients is precisely the target dimension $N$ of the map $f$.
+If the map is polynomial, that is $a=0$, then we already know from [19] that $f$ is unitarily equivalent to a juxtaposition of homogeneous maps.
+Since we are assuming $N=N_{f}$ in the theorem, none of these maps can be the constant.
+Thus since the map must be degree 2, it must be either equivalent to $H_{2}$ or a juxtaposition $\sqrt{1-t}H_{1}\oplus\sqrt{t}H_{2}$ for $t\in(0,1)$.
+We have therefore finished items (i) and (ii) of the theorem.
+In what follows, we will mostly consider the rational case $a>0$.
+Next we also notice that it is sufficient to consider the case $n=2$.
+This is because the right hand side of ˜47 is an expression that is a function of $z_{1}$, $\bar{z}_{1}$ and $\mathopen{}\mathclose{{\left\lvert z_{2}}}\right\rvert^{2}+\cdots+\mathopen{}\mathclose{{\left\lvert z_{n}}}\right\rvert^{2}$.
+Therefore, if we find the normal form in the case $n=2$, the larger dimensions will follow by replacing $\mathopen{}\mathclose{{\left\lvert z_{2}}}\right\rvert^{2}$ with $\mathopen{}\mathclose{{\left\lvert z_{2}}}\right\rvert^{2}+\cdots+\mathopen{}\mathclose{{\left\lvert z_{n}}}\right\rvert^{2}$ in the expression for $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$.
+Let us then fix $n=2$ and write the matrix of coefficients of the right hand side of ˜47.
+We order the monomials as $1,z_{1},z_{1}^{2},z_{2},z_{1}z_{2},z_{2}^{2}$.
+With this ordering the matrix becomes
+(48) $\begin{bmatrix}(1-b)+Qr^{2}&(1-b)a&0&0&0&0\\ (1-b)a&(1-b)a^{2}+b-Q(1+r^{2})&ba&0&0&0\\ 0&ba&ba^{2}+Q&0&0&0\\ 0&0&0&b-Q(1+r^{2})&ba&0\\ 0&0&0&ba&ba^{2}+2Q&0\\ 0&0&0&0&0&Q\end{bmatrix}$.
+[…]
+Call this coefficient matrix $M$.
+Notice that we picked an ordering of the monomials so that the matrix is a direct sum of three matrices, $M=M_{1}\oplus M_{2}\oplus M_{3}$ for a $3\times 3$ $M_{1}$, a $2\times 2$ $M_{2}$, and $M_{3}=[Q]$.
+The three matrices $M_{1}$, $M_{2}$, and $M_{3}$ must be positive semidefinite.
+Before we move on, we notice that given a right hand side of ˜47 for $n=2$, we can find a $p(z)$ where all components of $p(z)$ that depend on $z_{2}$ are either of the form $L(z_{1})z_{2}$ for an affine $L(z_{1})$, or the last term $\sqrt{Q}\,z_{2}^{2}$.
+That means that if we find a presentation of a representative of the equivalence class of maps for $n=2$, to get the corresponding map for $n>2$ we simply replace the multiplication by $z_{2}$ with tensoring by $(z_{2},\ldots,z_{n})$ and we replace $z_{2}^{2}$ by the symmetrized second tensor power of $(z_{2},\ldots,z_{n})$, that is, by $H^{\prime}_{2}(z_{2},\ldots,z_{n})$.
+That is where those terms in ˜44 come from.
+From now on we therefore generally assume that $n=2$.
+The block $M_{3}$ is trivial as $Q>0$ and gives rise to the term $\sqrt{Q}\,z_{2}^{2}$.
+We will show that the matrix $M_{2}$ is always positive definite (nonsingular in fact) when $M_{1}$ is positive semidefinite, and so it will not be an issue.
+Proposition 6.3.
+For a degree 2 annulus map $f$, we have $Q\leq 1$ and $0<b\leq 1+r^{2}$. In particular, we have that $R\geq r^{2}$. Furthermore, if $f$ is not polynomial, then $Q<1$ and $b<1+r^{2}$ (so $R>r^{2}$).
+Proof.
+Clearly $b>0$.
+If we consider the first diagonal entries of $M_{1}$ and $M_{2}$, we find that they must be nonnegative, that is, $1-b+Qr^{2}\geq 0$ and $b-Q(1+r^{2})\geq 0$.
+Together this means that $Q(1+r^{2})\leq b\leq 1+Qr^{2}$.
+This implies that $Q\leq 1$ and $b\leq 1+r^{2}$.
+Suppose now that $Q=1$.
+Then $b=1+r^{2}$ and also the first diagonal entry of $M_{2}$ is now 0.
+Thus for the matrix to be positive semidefinite we require the off-diagonal entry $ba=0$.
+Since $b>0$, we find that $a=0$.
+Thus if $a>0$, we have that $Q<1$.
+Consequently $b<1+r^{2}$.
+∎
+For the rest of the argument we will mostly consider the matrix
+(49) $M_{1}=\begin{bmatrix}(1-b)+Qr^{2}&(1-b)a&0\\ (1-b)a&(1-b)a^{2}+b-Q(1+r^{2})&ba\\ 0&ba&ba^{2}+Q\end{bmatrix}$.
+By setting $z_{2}=0$, we have that $M_{1}$ is a matrix of coefficients corresponding to a proper annulus map with source dimension 1.
+Assuming that $f$ does not map into a subspace, the matrix $M_{1}$ cannot be zero (as not all components of $p$ can be divisible by $z_{2}$).
+If the rank of $M_{1}$ is 1, then we find an annulus map from dimension 1 to dimension 1.
+We also know that the numerator $p$ has to be of degree $2$.
+This map therefore has to be $e^{i\theta}z^{2}$, meaning $Q=1$, $a=0$, and $b=1+r^{2}$.
+As we are assuming that $a>0$, this is not a possibility.
+Therefore, $M_{1}$ must be of rank 2 or 3.
+Proposition 6.4.
+Suppose $p$ has linearly independent components and $a>0$. The target dimension $N$ of $f$ is equal to the embedding dimension $N_{f}$ if and only if $M_{1}$ is of rank 2.
+Proof.
+We are left with $M_{1}$ being rank 2 or 3.
+If $M_{1}$ is of rank 3 and hence of full rank, then we can subtract from $M_{1}$ a multiple of the matrix $A$ corresponding to $\mathopen{}\mathclose{{\left\lvert 1+az_{1}}}\right\rvert^{2}$ and $M_{1}-\epsilon A$ is still positive (semi)definite for small enough $\epsilon>0$.
+Pick $\epsilon$ large enough so that $M_{1}-\epsilon A$ is rank 2, which we decompose as a sum of outer products as usual.
+Thus we find a decomposition $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ with linearly independent components, where one of the components of $p$ is $\sqrt{\epsilon}(1+az_{1})$.
+But that means that the corresponding $f=\frac{p}{q}$ has a constant term, meaning that the embedding dimension for $f$ is strictly smaller than $N$.
+Conversely, suppose $M_{1}$ is rank 2 and the embedding dimension is strictly smaller than $N$.
+This means that after a unitary, we can make one of the components of $f$ be a constant, meaning that $\mathopen{}\mathclose{{\left\lVert f(z_{1},0,\ldots,0)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z_{1})}}\right\rvert^{2}+c^{2}$, as only the $M_{1}$ block contributes terms and the $\mathopen{}\mathclose{{\left\lvert f_{1}(z_{1})}}\right\rvert^{2}$ comes from the second eigenvalue of $M_{1}$.
+Again, this means that $f_{1}$ is a proper map of annuli in the complex plane and hence just a multiple of $z^{2}$ and this leads to a contradiction if $a>0$ as before.
+∎
+As we are assuming that $M_{1}$ is singular, we have the equation
+(50) $0=\det(M_{1})=-r^{2}(r^{2}+1)Q^{3}\\ +(-a^{2}br^{4}-2a^{2}br^{2}+2br^{2}+a^{2}r^{2}-r^{2}+b-1)Q^{2}+(1-a)(1+a)(1-b)b(1-ar)(ar+1)Q.$
+This equation is divisible by $Q$.
+The root at $Q=0$ is not interesting as we are interested in $Q>0$, so we can discard it and we are left with a quadratic equation.
+The two solutions to this quadratic equation are written out in ˜46.
+As a cubic in $Q$, $\det(M_{1})$ has negative cubic term.
+It therefore goes to negative infinity as $Q\to\infty$.
+It is now good to consider the cases $b<1$ ($R>r$), $b=1$ ($R=r$), and $b>1$ ($R<r$) separately. We start with $b<1$. In this case, 1b>01-b>0. We note that
+(51) $\mathopen{}\mathclose{{\left.\frac{\partial}{\partial Q}}}\right|_{Q=0}\det(M_{1})=(1-a)(1+a)(1-b)b(1-ar)(ar+1)>0.$
+This means that as a function of $Q$, $\det(M_{1})$ must have a negative root $Q_{-}$, the aforementioned root at 0, and a positive root $Q_{+}$.
+By looking at the matrix $M_{1}$ when $a=0$, we find that it has 3 positive eigenvalues when $Q>0$ is small enough.
+As the sign of an eigenvalue can only change when the determinant goes through 0, we have that $M_{1}$ is positive semidefinite for all $0\leq Q\leq Q_{+}$.
+For similar continuity reasons the same is true for all $0<a<1$.
+As $\det(M_{1})$ is negative for all $Q>Q_{+}$, the matrix $M_{1}$ cannot be positive semidefinite for $Q>Q_{+}$.
+Next we still need to handle what happens at $Q=1$.
+While we have shown that for a map to exist, we must have $Q<1$, what we want to show is that $Q_{+}<1$ and hence a map always exists.
+For this we plug in $Q=1$ into $\det(M_{1})$.
+(52) $\det(M_{1})|_{Q=1}=-(a^{2}b+1)r^{4}-(a^{4}b^{2}-a^{2}b^{2}-a^{4}b+3a^{2}b-2b-a^{2}+2)r^{2}-(1-b)(a^{2}b+1-b).$
+That is, we have a polynomial in $r$.
+It is clear that for the considered values of $a\in(0,1)$ and $b\in(0,1)$ the coefficient of $r^{4}$ and the constant are strictly negative.
+The coefficient of $r^{2}$ is also negative in this range which is a simple calculus exercise as the expression has a strict minimum on the unit square at $a=0,b=1$.
+By Descartes’ rule of signs, there are no positive solutions $r$ for $\det(M_{1})|_{Q=1}=0$.
+Therefore $\det(M_{1})|_{Q=1}$ is always strictly negative for the allowed $a$, $b$, and $r$, and hence $Q_{+}<1$.
+We need to next show that for the range of $Q$ where $\det(M_{1})$ is nonnegative, we have that $\det(M_{2})>0$. We have
+(53) $\det(M_{2})=(-2(r^{2}+1)Q+(2-a^{2}r^{2}-a^{2})b)Q.$
+This function has two roots, one at the origin and one positive, call it $Q_{2,+}$.
+It is easy to see that at $a=0$ the matrix is positive semidefinite and since $\mathopen{}\mathclose{{\left.\frac{\partial}{\partial Q}}}\right|_{Q=0}\det(M_{2})>0$ we find that $M_{2}$ is positive semidefinite between the two roots.
+We plug $Q=Q_{2,+}$ into $\det(M_{1})$ to find
+(54) $\frac{a^{2}b^{2}(2-a^{2}r^{2}-a^{2})\bigl((a^{2}r^{6}-a^{2}r^{2}-2r^{2}+2)b+(2a^{2}r^{4}-2r^{4}+2a^{2}r^{2}-2)\bigr)}{8(r^{2}+1)^{2}}.$
+The first two terms of the numerator are clearly strictly positive in our allowed range.
+It is also not hard to see by basic calculus that the third term of the numerator is strictly negative (it grows as $b$ grows and is still negative at $b=1$).
+In other words, we have shown that at $Q=Q_{+}$, $M_{2}$ is still positive definite.
+Hence, we have a coefficient matrix $M$ that is positive semidefinite and writing the blocks of $M$ as sums of hermitian squares obtains the form in the theorem.
+As $M_{2}$ is nonsingular and $M_{1}$ is rank 2, it is not hard to check that the matrix $M$ is of full rank minus 1 for any $n\geq 2$.
+Hence we are finished with $b<1$ and the item (v) of the theorem.
+Let us now consider $b=1$. When $b=1$, we find that $\det(M_{1})$ simplifies greatly to
+(55) $\bigl(r^{2}(1-a^{2}(1+r^{2}))-Qr^{2}(1+r^{2})\bigr)Q^{2}.$
+There are two roots at $Q=0$ and one other root $Q_{0}=\frac{1}{1+r^{2}}-a^{2}$.
+If the root is negative then the determinant is negative for positive $Q$, hence we only have a map if the root is positive.
+It is positive if and only if $\frac{1}{\sqrt{1+r^{2}}}>a$, which gives the restriction on $a$ in item (iv) of the theorem.
+It is also clear that $Q_{0}<1$.
+Moreover, plugging $Q=Q_{0}$ into $\det(M_{2})$ we obtain
+(56) $a^{2}\bigl(1-a^{2}(1+r^{2})\bigr).$
+Thus $\det(M_{2})$ is strictly positive as are the two diagonal terms, and so the matrix $M_{2}$ is positive definite. Item (iv) of the theorem now follows by expanding $M$ in the same way as before.
+Finally, we consider $1<b<1+r^{2}$.
+The coefficient of $Q$ in $\det(M_{1})$ is (1a)(1+a)(1b)b(1ar)(ar+1)(1-a)(1+a)(1-b)b(1-ar)(ar+1) and so we see that the root at $Q=0$ is always a simple root.
+If we plug in some sample values, we find that it is possible to obtain positive roots.
+Hence the two roots of $\det(M_{1})=0$ are always positive or possibly not real.
+When $a=0$, the matrix $M_{1}$ has two positive and one negative eigenvalues for negative $Q$ and has roots $Q_{1}=\frac{b-1}{r^{2}}$ and $Q_{2}=\frac{b}{1+r^{2}}$, which satisfy $0<Q_{1}<Q_{2}<1$.
+It is not hard to see that $M_{1}|_{Q=1}$ is not positive semidefinite.
+Otherwise the center term would be nonnegative: $(1-b)a^{2}+b-(1+r^{2})\geq 0$.
+Together with $b<1+r^{2}$, we would obtain $(1-b)a^{2}>0$, which would contradict $1<b$.
+Therefore, as $a$ grows the two roots must stay less than $1$ until we either reach $a=1$ or they disappear if the discriminant becomes zero.
+In this range, as the matrix $M_{1}$ is positive semidefinite between the two roots at $a=0$ it is also positive semidefinite throughout.
+To find if and where the two roots disappear, we consider the discriminant of the quadratic $\det(M_{1})/Q=0$:
+(57) $a^{4}r^{4}(br^{2}+1)^{2}-2a^{2}r^{2}(br^{2}(r^{2}+b-1)+1+r^{2}-b)+(1+r^{2}-b)^{2}.$
+This expression is strictly positive when $a=0$ and at $a=1$ it is $(1-b(1-r^{4}))^{2}\geq 0$.
+The expression can be considered a quadratic in $a^{2}$ and it is concave up.
+To show that it always has two roots between 0 and $1$, it is better to substitute $b=1+sr^{2}$ so that we have a parameter $s\in(0,1)$ just like $a$ and $r$.
+The minimum of this polynomial is at
+(58) $a^{2}=\frac{r^{4}s^{2}+r^{4}s+r^{2}s-s+r^{2}+1}{(r^{4}s+r^{2}+1)^{2}}.$
+It is a calculus exercise that for $s,r$ in $(0,1)$, the $a$ where this minimum occurs is always in $(0,1)$.
+We plug the location of this minimum into the discriminant ˜57 (with $b=1+sr^{2}$) to obtain
+(59) $\frac{-4r^{6}(r^{2}+1)s(r^{2}s+1)(r^{2}s-s+1)}{(r^{4}s+r^{2}+1)^{2}}.$
+It is not hard to see that this minimum is then always negative.
+Putting this all together means that the discriminant ˜57 always has two positive roots, both of these roots are within $(0,1]$, with the smaller one strictly less than $1$.
+Let $a_{1}$ and $a_{2}$ denote the two roots where $a_{1}$ is the smaller one.
+The smaller root is given as the right hand side in ˜45.
+For $0<a\leq a_{1}$, we therefore obtain two roots $Q_{1}$ and $Q_{2}$ in $(0,1)$ to $\det(M_{1})=0$ and between them the matrix $M_{1}$ is positive semidefinite, which can be checked by plugging in a test point between the two roots and noting that by continuity the entire connected region given by $Q_{1}<Q<Q_{2}$, $0<a<a_{1}$, $s\in(0,1)$, where the determinant is strictly positive, we must get that the matrix is positive definite.
+Thus it is semidefinite when we allow nonstrict inequalities.
+Similarly we check the regions $Q_{1}<Q<Q_{2}$, $a_{1}<a<1$, $s\in(0,1)$ (there are two such regions as $a_{1}=1$ if $b=\frac{1}{1-r^{4}}$ as we have seen above).
+In both of these regions the matrix $M_{1}$ has a negative eigenvalue.
+We are done with the matrix $M_{1}$.
+It remains to show that $M_{2}$ is positive definite if the parameters are within the stated limits.
+This is not hard to see, if we delete the first row and first column of $M_{1}$ we get a positive semidefinite matrix $M_{1}^{\prime}$.
+Then $M_{2}$ can be obtained from $M_{1}^{\prime}$ by adding strictly positive values to both diagonal entries, hence $M_{2}$ is in fact positive definite.
+Item number (iii) now follows.
 7. Other remarks on proper maps of annuli
@@
 Since the automorphism group of the ball is transitive, every spherical equivalence class of proper maps of balls contains a map taking 0 to 0.
-The case of the complement of the closed unit ball was studied in [helal-2025-proper].
+The case of the complement of the closed unit ball was studied in [19].
 In that case, for $n\geq 2$, the proper maps are polynomial sphere maps whose norm goes to infinity at infinity.
@@
 For homogeneous maps as above we have $r^{d}\leq R$, with $H_{d}$ itself achieving the equality in this bound.
-Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general.
-If one also requires that $N$ is the embedding dimension ($N=N_{f}$), then more can be said.
-In this case, if $d=1$, then $R=r$.
-For higher degree, one could use juxtaposition to get a range.
-For example, if we juxtapose $H_{1}$ and $H_{d}$, then via the calculation in the introduction, we have that for $t\in(0,1)$, the map $\sqrt{1-t}\,H_{1}\oplus\sqrt{t}\,H_{d}$ has $R=\sqrt{(1-t)r^{2}+tr^{2d}}$, and so any $R$ in the range between $r^{d}$ and $r$ is achieved by these maps.
-So the question could be modified for the case when $N=N_{f}$ and degree $d>1$, to ask if $R$ equal to or bigger than $r$ is possible.
+Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems to be if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general.
+In Section˜6, we show that all degree 2 maps satisfy the bound $r^{2}\leq R$, giving further support for this conjecture.
+Moreover, we construct degree 2 maps with $r^{2}<R<1$ such that the target dimension $N$ is the embedding dimension.
+That means that there is no simple upper bound on $R$ if the degree $d$ is larger than 1.
 Finally, we remark that given our results and our examples, it is plausible to conjecture a degree bound in terms of the dimensions.